### v1

# D-NPC: Dynamic Neural Point Clouds for Non-Rigid View Synthesis from Monocular Video

Abstract
Dynamic reconstruction and spatiotemporal novel-view synthesis of non-rigidly deforming scenes recently gained increased attention. While existing work achieves impressive quality and performance on multi-view or teleporting camera setups, most methods fail to efficiently and faithfully recover motion and appearance from casual monocular captures. This paper contributes to the field by introducing a new method for dynamic novel view synthesis from monocular video, such as casual smartphone captures. Our approach represents the scene as a dynamic neural point cloud, an implicit time-conditioned point distribution that encodes local geometry and appearance in separate hash-encoded neural feature grids for static and dynamic regions. By sampling a discrete point cloud from our model, we can efficiently render high-quality novel views using a fast differentiable rasterizer and neural rendering network. Similar to recent work, we leverage advances in neural scene analysis by incorporating data-driven priors like monocular depth estimation and object segmentation to resolve motion and depth ambiguities originating from the monocular captures. In addition to guiding the optimization process, we show that these priors can be exploited to explicitly initialize our scene representation to drastically improve optimization speed and final image quality. As evidenced by our experimental evaluation, our dynamic point cloud model not only enables fast optimization and real-time frame rates for interactive applications, but also achieves competitive image quality on monocular benchmark sequences. Our project page is available at https://moritzkappel.github.io/projects/dnpc.
[…]
figure We introduce Dynamic Neural Point Clouds for dynamic view synthesis from monocular videos. Our method enables fast optimization, real-time frame rates, and competitive image quality.
1 Introduction
Synthesizing novel views from a sparse set of input images is a fundamental challenge in computer vision. Using recent advances in 3D neural scene reconstruction, it becomes possible to create photorealistic renderings from arbitrary viewpoints, effectively adding interactive six degrees-of-freedom camera support. In the context of general, non-rigidly deforming dynamic scenes, the capability of interactively rendering novel views for every point in time enables numerous applications, including video stabilization [28], immersive playback for social media [2], or virtual reality [57]. However, performing faithful reconstruction from just a single monocular video (e.g., a casual smartphone recording) remains a computationally expensive and challenging problem.
While the introduction of Neural Radiance Fields [31] and 3D Gaussian Splatting [19] brought about significant advances in static scene reconstruction, dynamic extensions still lag behind in terms of quality and robustness due to motion and depth ambiguities in the temporal domain. Those ambiguities are usually resolved using expensive large-scale multi view setups [24] or monocularized data (i.e., teleporting camera setups) [18], which maximize the ratio between camera and scene motion to improve temporal consistency. However, recent investigations [15] reveal that most sparsely regularized methods fail to faithfully recover consistent motion and geometry from casual smartphone captures due to the lack of effective multi-view signal. Multiple recent radiance field approaches address this problem by incorporating strong priors like monocular depth, optical flow, and semantic segmentation for the joint reconstruction of geometry, 4D scene flow fields [25, 26], and even camera parameters [27] from monocular video. While current monocular methods achieve high-quality spatiotemporally consistent view synthesis, they are often limited in the expressiveness of novel camera views, and by their high computational demand.
To address the above-mentioned challenges, this paper introduces Dynamic Neural Point Clouds (D-NPC), a new system for monocular dynamic view synthesis based on implicit point distributions and fast hash-encoded feature grids. Given a monocular input video and the corresponding structure-from-motion calibration, our method optimizes a 4D representation of the depicted scene, which enables spatio-temporal novel-view synthesis, e.g., the infamous bullet-time effect. Following the concept of implicit neural point clouds [17], we track expected scene density in a sparse voxel grid, which is used to extract explicit pose-dependent point clouds. Combined with a fast multi-resolution hash-encoded feature grid [33], this scene representation is able to efficiently render high-fidelity novel views using a differentiable rasterizer and image-to-image translation network. Based on the concepts of space-time hash encoding [51], we lift the representation to the temporal domain by replacing the single feature grid with a 3D and 4D hash encoding for the static background and dynamic foreground regions, respectively. We further extend the point distribution field to robustly track the per-voxel ratio of static and dynamic scene content, effectively enabling 3D dynamics segmentation to significantly reduce the required compute during optimization and inference.
To guide the severely under-constrained monocular reconstruction, we leverage per-frame data-driven priors, namely optical flow, monocular depth, and foreground segmentation masks, which we extract and preprocess with minimal overhead. We demonstrate that our temporal extension can take advantage of the rasterization-based forward rendering of implicit neural point clouds, which enables efficient foreground-background decomposition and image-space regularization. Furthermore, the scene-level sampling approach also provides a seamless way of integrating data-driven priors for model initialization, resulting in faster and better convergence during optimization. As shown in D-NPC: Dynamic Neural Point Clouds for Non-Rigid View Synthesis from Monocular Video, D-NPC achieves competitive image quality while significantly reducing optimization time, and allowing for interactive rendering in a graphical user interface.
Overall, we summarize our main technical contributions as follows:
  • We propose a new, efficient dynamic scene representation for non-rigid scene reconstruction and novel-view synthesis from a single monocular input video. D-NPC is the first adaptation of implicit neural point clouds for dynamic view synthesis.
  • We demonstrate that our model can elegantly leverage additional priors like monocular depth and object segmentation masks to explicitly initialize the sampling distribution field, which drastically improves optimization and inference speed.
Our method enables fast optimization in as few as 0.27 GPU hours and interactive frame rates above 60 FPS during inference, while achieving competitive image quality on common benchmark datasets.
2 Related Work
Rendering novel viewpoints of a recorded scene is a longstanding and well-studied problem. Early work on image-based rendering uses morphing to generate in-between views from image collections without explicit 3D reconstruction [9, 42]. Other approaches apply implicit or explicit 3D scene representations like light fields [16] or layered depth images [43] to generate high-quality novel views. Recently, Neural Radiance Fields (NeRFs) [31] and 3D Gaussian Splatting (3D-GS) [19] inspired a multitude of follow-up research, including extensions for dynamic scenes. We next briefly summarize neural representations for static scenes, followed by dynamic view synthesis.
2.1 Static View Synthesis
NeRFs [31] implicitly represent the scene geometry and appearance using a Multi-Layer Perceptron (MLP), yielding an unprecedented image quality for novel views at the cost of slow optimization and rendering due to the large number of network queries required. While some follow-ups propose qualitative improvements for, e.g., anti-aliasing and unbounded scenes [3, 4, 5], another branch of methods focuses on acceleration using explicit scene representations like sparse [13] or decomposed [8] voxel grids. Along the same lines, Müller et al. [33] propose an efficient grid-based scene representation implemented as a multi-resolution hash-grid, enabling high-quality reconstruction in only a few minutes. All the above volume rendering approaches apply backward rendering, where the scene is frequently sampled along the camera rays. In contrast, forward rendering approaches, explicitly model scene geometry using geometric primitives, allowing for fast rasterization-based rendering. The popular 3D-GS by Kerbl et al. [19] represents the scene as a set of 3D Gaussians for very fast rendering at a high image quality, inspiring many extensions [22, 60]. Analogously, point-based rendering approaches [1, 21, 39] use a fixed initial point cloud combined with a neural image translation network for hole filling. The recent INPC approach by Hahlbohm et al. [17] combines the benefits of forward and backward rendering, modelling the scene as a point probability and hash-encoded appearance field for high-quality rendering at interactive frame rates. Our D-NPC method leverages these benefits for fast dynamic view synthesis from monocular video. A more detailed discussion of neural rendering techniques can be found in the report by Tewari et al. [48].
2.2 Dynamic View Synthesis
Most NeRF-based extensions use a time-conditioned MLP [24, 25, 36] or additional deformation field [11, 34, 35, 37] to represent scene dynamics. Following acceleration models for static reconstruction, similar models based on factorized [2, 7, 12, 52] and hash-encoded [18, 45, 51] feature grids were introduced for dynamic view synthesis. D2NeRF [56] combined two models for foreground-background separation from a single video. Likewise, Wang et al. [51] decompose the scene in a fast 3D and 4D hash-grid representation. We adapt concepts from both approaches, integrating foreground/background handling in both our sampling and appearance modelling, enabling our method to individually render dynamic and static scene content while improving performance. Similar to radiance fields, recent approaches opt to extend 3D-GS for non-rigid scenes, tracking the 3D motion of Gaussians over time [29, 55] while retaining fast rendering. However, as recently revealed by Gao et al. [15], these methods require multi-view or monocularized [18] capturing setups to exploit observations from multiple views. On the other hand, multiple methods focus on dynamic view synthesis from a single monocular input video, integrating additional priors like dense scene flow estimation to resolve motion and depth ambiguities [25, 26, 27]. RoDynRF [27] further takes the challenging scenario to the extreme, additionally optimizing camera poses from the unposed input video. However, these MLP-based monocular approaches rely on backward rendering, thus being limited in optimization speed and final rendering frame rate. This paper presents a point-based forward rendering approach for monocular view synthesis, which leverages additional priors to initialize and guide the reconstruction, enabling high-quality rendering at real-time frame rates. More details on the state of the art in dynamic reconstruction can be found in the survey by Yunus et al. [61].
3 Method
Figure 1: Method overview. Given a monocular RGB video and priors extracted from out-of-the-box estimators, we initialize and optimize a dynamic implicit neural point cloud, consisting of a spatiotemporal point position distribution and two feature grids for static and dynamic scene content. By sampling an explicit point cloud for a discrete timestamp, our model can synthesize novel views, including foreground/background separation, using a differentiable rasterizer and neural renderer.
We introduce Dynamic Neural Point Clouds (D-NPC), a method for dynamic novel-view synthesis from monocular video. Given a single monocular video consisting of $N$ images $I=(i_{1},...,i_{N})$ and their normalized timestamps $T=(t_{1},...,t_{N})\in[0,1]^{N}$, we reconstruct a dynamic implicit point cloud representation [17] of the depicted dynamic scene. Based on its static counterpart, the core of our D-NPC model consists of a spatiotemporal point probability ($\mathcal{P}$) and appearance feature descriptor ($\mathcal{A}$) field, which can efficiently be rendered using a differentiable rasterizer and neural rendering network. A full overview of our method is given in Fig. 1. In the following, we describe the individual components and optimization process in detail.
3.1 Data Preprocessing
The reconstruction of non-rigidly deforming scenes from a single moving camera is a highly ill-posed problem. To regularize the reconstruction from monocular video, previous approaches [15, 25, 27] apply estimates from several image processing applications, replacing the multi-view consistency signal with learned, data-driven scene analysis. Following this idea, our method starts with data preprocessing to extract high-quality priors, which we use to initialize and guide our model in the later stages. Despite the recent emergence of consistent yet expensive scene calibration [44], video depth estimation [30] and motion tracking [53] estimators, one of our method’s main objectives is speed and simplicity. Thus, we focus on real-time out-of-the-box solutions to maintain a fast total reconstruction time (including prior computation).
First, we start by applying sparse structure-from-motion (SfM) using COLMAP [40, 41] to estimate camera intrinsics and per-frame poses (extrinsics) $E=(e_{1},...,e_{N})$. As a side product, we extract the sparse, globally-aligned SfM point cloud. We further use RAFT [47] to estimate dense optical flow in the forward and backward directions, and estimate per frame (dis-)occlusion masks based on forward-backward consistency. To reduce the depth ambiguity of dynamic scene content, we employ the fast monocular depth estimator DepthAnything [58], which provides relative disparity maps for every video frame. Instead of using a relative depth loss [25], we align the predicted disparity with the global SfM point cloud by estimating a scale and shift via RANSAC-based least squares and temporal smoothing. This way, we obtain globally-aligned depth which can explicitly be integrated into our model. Lastly, we obtain binary segmentation masks $\omega$ of moving foreground object(s) to mask out dynamic scene content in our reconstruction and COLMAP SfM feature extraction. Similar to monocular depth, automatic motion segmentation is highly ill-posed, even in the presence of semantic segmentation masks. Thus, we use the semi-automatic video object segmentation system Cutie [10], which is guided by minimal user input in the form of manual keypoint selection. In line with our demands on simplicity, Cutie generates masks of sufficient quality in less than a minute, requiring as few as two mouse clicks, which in the future could be integrated into the data recording setup on mobile devices.
3.2 Dynamic Neural Point Cloud Representation
Our method represents the 4D scene as a Dynamic Neural Point Cloud, the first temporal extension of implicit neural point clouds [17]. The primary idea of implicit point clouds is to replace the persistent point cloud $(p,f)$, consisting of a set of 3D positions $p$ and corresponding color or feature descriptor vectors $f$, with a sampling distribution $\mathcal{P}$ and a neural feature grid $\mathcal{A}$. By repeatedly sampling and rendering an explicit point cloud $((p,\mathcal{A}(p)),p\sim\mathcal{P})$, it is possible to iteratively update and refine the spatial distribution, while maintaining all benefits of point-based rendering. However, applying this concept to dynamic scenes is not straightforward, as the direct integration of a fourth dimension is memory-intensive and does not naturally exploit any temporal coherency.
3.2.1 Temporal Point Probability Field
The probability field $\mathcal{P}$ models the likelihood for spatial point sampling by tracking the estimated local scene opacity $\sigma\in[0,1]$ over the course of optimization. To facilitate accurate and efficient sampling over a dynamic sequence, we divide our probability field $\mathcal{P}=(\mathcal{P}_{s},\mathcal{P}_{d})$ into two separate instances for static and dynamic regions, respectively. While the static 3D field $\mathcal{P}_{s}$ continues to track local opacity $\sigma_{s}$ assumed to be consistent over all points in time, the 4D dynamic field $\mathcal{P}_{d}$ stores additional values $\sigma_{d}(t)$ as a function of time. To enable a clean separation of foreground and background, $\mathcal{P}_{d}$ not only models the expected local opacity, but additionally keeps track of the estimated local proportion $\beta(t)$ between static and dynamic substance, as explained later on. In conjunction, our probability fields can sample explicit, time-aware point positions, while pre-marking individual points as either dynamic or static to reduce compute complexity. Given the natural imprecisions and relatively low resolution of monocular reconstruction, we find that the octree-based data structure suggested by Hahlbohm et al. [17] does not yield significant benefits due to the absence of fine geometric details. Instead, we use a fixed-resolution sparse voxel grid representation for the static and dynamic fields.
Voxel Grid Initialization.
Despite using a sparse data structure, starting from a fully occupied grid (i.e., all cells are sampled equally for all timestamps) is exceedingly inefficient regarding both memory consumption and model convergence. Thus, we make use of the data-driven priors described in Sec. 3.1 to find a sparse initial sampling distribution before optimization. First, we initialize the voxel grid bounds as the minimal cuboid enclosing all SfM points after performing outlier filtering based on the 95-percentile, containing a fixed total of $128^{3}$ and $N\cdot 128^{3}$ voxels for $\mathcal{P}_{s}$ and $\mathcal{P}_{d}$ respectively. Then, for every timestamp $t_{i}\in T$, we use the corresponding training camera matrix $e_{i}$ to project a set of random points in each voxel to the image plane, and gather the minimal relative distance to the estimated monocular depth $d_{i}$, and the maximum over the binary foreground segmentation mask $\omega_{i}$. We set the initial opacity-based sampling probability $\sigma_{d}(t_{i})$ of each cell to the relative deviation from the monocular depth, using the maximum over all timestamps for the static field ($\sigma_{s}=\max(\{\sigma_{d}(t)\>|\>t\in T\})$). Additionally, we carve the temporal slices of $\mathcal{P}_{d}$ using the dynamic segmentation. By thresholding the sampling probabilities, we eliminate up to 95% of voxels, resulting in an efficiently represented distribution constituting a proper local minimum for optimization.
Explicit Point Cloud Sampling.
For an arbitrary camera view $e$ and timestamp $t$, we sample a set $P\in\mathbb{R}^{k\times 3}$ of $k$ explicit point positions ($P\sim\mathcal{P}\>|\>(e,t)$) for rendering a novel view. To this end, we first take the union of all voxels in the static field and the dynamic field at time $t$, taking the maximum expected opacity of voxels existing in both models. Then, we eliminate voxels based on the camera frustum by projecting each voxel center to the image plane using $e$. To obtain the final point positions $p$, we interpret the per-voxel opacity as a discrete PDF, using discrete multinomial sampling (with replacement) to propose a list of voxels, in each of which we place a single uniformly sampled 3D position. This way, more points are sampled in areas with a high expected scene density, which can iteratively be refined during optimization. For the resulting set of point positions, we further trace a binary mask $m_{p}$ indicating if a point was sampled from static or dynamic scene regions, enabling efficient model evaluation by omitting excessive queries of the dynamic appearance model.
Voxel Grid Update.
We update the sampling distribution field by tracking the blending weights assigned to every sampled explicit point during rasterization. We first apply a constant decay factor to the values of every unique voxel proposed by the multinomial sampling, and then assign it the maximum over the current value and the blending weights of all points sampled from the specific cell. For samples from dynamic voxels of $\mathcal{P}_{d}$, we apply the same procedure to the tracked static-to-dynamic ratios $\delta$ using per-point densities described in Sec. 3.2.2. Afterwards, we prune voxels from the sampling distribution by threshold: For $\mathcal{P}_{d}$, we remove cells with a low $\delta$-value, as it either describes empty space or is well handled by the static model. We further remove voxels $\mathcal{P}_{s}$ from $\mathcal{P}_{d}$ whose expected opacity $\sigma$ is low in the static 3D field and for all remaining timestamps in the dynamic field. This way, we progressively reduce memory consumption and refine importance sampling.
3.2.2 Dynamic Appearance Field
In a conventional point cloud, every point is assigned a fixed feature vector describing its appearance. As implicit neural point clouds [17] repeatedly extract new sets of point positions, they use a fast multi-resolution hash-encoded feature grid [33] $\mathcal{A}$ to model the appearance over the entire spatial sampling distribution. To represent time-conditioned appearance for dynamic scenes, we again use a static 3D and dynamic 4D hash-encoded neural feature grid $\mathcal{A}=(\mathcal{A}_{s},\mathcal{A}_{d})$. As previously described by Wang et al. [51], using an ensemble of 3D and 4D hash-grids instead of a single 4D representation effectively improves the capabilities of the dynamic model by reducing hash collisions. To model the static and dynamic scene appearance at a spatial input position $p$ at time $t$, the appearance fields yield a density $\delta\in\mathbb{R}_{0}^{+}$ and an $n_{f}$-dimensional feature vector $f\in[0,1]^{n_{f}}$:
$\delta_{s},f_{s}=\mathcal{A}_{s}(p),\quad\delta_{d},f_{d}=\mathcal{A}_{d}(p,t).$ (1)
Note that we only have to query $\mathcal{A}_{d}$ for points indicated as dynamic by the binary segmentation mask $m$, which significantly reduces compute and memory. Following the formulation of D2NeRF [56], we then calculate the combined per-point density $\delta_{c}$ as the sum of static and dynamic densities, and obtain the final feature descriptor $f_{c}$ via density-weighted interpolation:
$\delta_{c}=\begin{cases}\delta_{s}+\delta_{d}&\text{if}\quad m(p)\\ \delta_{s}&\text{else}\end{cases},\qquad f_{c}=\begin{cases}\frac{\delta_{s}f_{s}+\delta_{d}f_{d}}{\delta_{c}}&\text{if}\quad m(p)\\ f_{s}&\text{else}\end{cases}.$ (2)
For rendering and probability field updates, we then transform density values $\delta$ into opacity $\sigma$, and calculate the dynamic ratio $\beta$ via:
$\sigma_{(s/d/c)}=1-e^{-\delta_{(s/d/c)}},\quad\beta=\frac{\delta_{d}}{\delta_{c}}.$ (3)
3.3 Neural Point Cloud Rendering
Given the time-dependent neural point cloud for an input view consisting of point positions $p$, point opacity $\sigma$, and appearance features $f$, we implement a fast differentiable point rasterizer to generate alpha ($\alpha$), depth ($d$), and 2D image feature maps $f_{\text{img}}$. Naturally, the rasterized images still contain holes, a common characteristic of classical point rendering approaches. Thus, we use a neural renderer, implemented as a light-weight convolutional neural network with a U-Net [38] architecture to fill the remaining holes, and convert the neural feature map $f_{\text{img}}$ to a human-interpretable RGB image $c$. In practice, we apply the rasterization and rendering components to the static, dynamic, and combined point cloud features separately, resulting in individual depth $d_{s,d,c}$ and color $c_{s,d,c}$ images that can be used for supervision and regularization. As the resolution of dynamic view synthesis benchmarks is usually low, we do not use any splatting or multi-resolution inputs in the rasterizer or neural renderer, which further increases performance.
3.4 Optimization
During training, we iteratively sample an explicit point cloud (Sec. 3.2.1) for a random training view and its associated timestamp, extract per-point appearance features (Sec. 3.2.2), render the resulting images (Sec. 3.3), update the sampling distribution (Sec. 3.2.1), and optimize model parameters using the following objective function:
$\mathcal{L}=\lambda_{\text{photo}}\mathcal{L}_{\text{photo}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{seg}}\mathcal{L}_{\text{seg}}+\lambda_{\text{dist}}\mathcal{L}_{\text{dist}},$ (4)
with hyperparameters $\lambda_{\text{photo}}$, $\lambda_{\text{depth}}$, $\lambda_{\text{seg}}$ and $\lambda_{\text{dist}}$. We use the same set of hyperparameters across all our experiments. The loss $\mathcal{L}_{\text{photo}}$ enforces photometric consistency between the rendered ($c$) and GT RGB image. Here, we use a weighted combination of pixel-level Cauchy loss [6], structural dissimilarity [54], and perceptual LPIPS [62] to maximize image quality. Our depth regularization loss $\mathcal{L}_{\text{depth}}$ penalizes the difference between the rasterized dynamic depth $d_{d}$ and the aligned monocular depth prior $d_{\text{gt}}$ in the binary foreground mask $\omega$:
$\mathcal{L}_{\text{depth}}=\frac{1}{\lvert\omega\rvert}\sum_{i=0}^{\lvert\omega\rvert}\omega_{i}\lvert d_{d,i}-d_{\text{gt},i}\rvert.$ (5)
The segmentation regularizer $\mathcal{L}_{seg}$ encourages a clean separation between the static and dynamic models via skewed binary entropy [56] on the dynamic ratios $\beta$:
$\mathcal{L}_{\text{seg}}=-\beta^{k}\log_{2}(\beta^{k})-(1-\beta^{k})\log_{2}(1-\beta^{k}),$ (6)
where $k$ is the skewing bias. As our last regularizer $\mathcal{L}_{dist}$, we adapt the distortion loss of Mip-NeRF360 [4, 46] to encourage thin surfaces, as it is not feasible to reliably reconstruct the time-dependent volume of a deforming object from a single monocular observation. For a detailed description of our training and loss schedule, please see our supplemental material (Sec. A.2).
4 Experimental Evaluation
We perform quantitative (Sec. 4.1) and qualitative (Sec. 4.2) evaluation and comparisons against state-of-the-art methods on two benchmark datasets: First, we evaluate our approach on the NVIDIA dynamic scenes dataset [59] (MIT license), featuring forward-facing scenes captured by a multi-view rig, monocularized by using a single frame per time to get a smooth camera trajectory. We extract training data using the procedure of RoDynRF [27], resulting in $12$ training and test views. For comparisons against DynIBaR [26] using their train/test split containing $\sim 150$ training views per scene, please see our supplemental material. Additionally, we evaluate our approach on the iPhone dataset [15] (Apache 2.0 license) which comprises challenging sequences with complex camera and scene motion trajectories, fine texture details, and (self-)occlusions. Furthermore, we provide an ablation study assessing the influence of individual pipeline components in Sec. 4.3.
4.1 Quantitative Results
Table 1: Timings on the NVIDIA dataset. We report training time and inference frame rates on the NVIDIA dataset with an image resolution of $480{\times}270$ pixels. Values for methods marked with $\dagger$ were gathered from the literature.
Method GPU Hours [h] $\downarrow$ Frame Rate [FPS] $\uparrow$
RoDynRF [27] 28 ${<}1$
HyperNeRF [35] 64 ${<}1$
DynamicNeRF [14] 74 ${<}0.1$
NSFF [25] 223 ${<}1$
DynIBaR [26] 320 ${<}1$
4D-GS [55] 1.1 45
Ours 0.27 70
Table 2: Quantitative comparisons on the NVIDIA and iPhone datasets against recent dynamic view synthesis approaches. Our method achieves the best average scores in terms of perceptual image quality. The three best results are highlighted in green in descending order of saturation.
NVIDIA Average
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
T-NeRF [37] 18.33 0.436 0.511
D-NeRF [37] 21.49 0.629 0.320
HyperNeRF [35] 17.60 0.377 0.502
NR-NeRF [49] 19.69 0.510 0.434
NSFF [25] 24.33 0.751 0.257
DynamicNeRF [14] 26.10 0.837 0.142
RoDynRF [27] 25.89 0.854 0.110
TiNeuVox [11] 19.74 0.501 0.412
4D-GS [55] 22.89 0.731 0.237
Ours 25.64 0.845 0.109
iPhone Average
Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
HyperNeRF+B+D+S [35] 16.81 0.569 0.332
Nerfies+B+D+S [34] 16.45 0.570 0.339
T-NeRF [15] 15.70 0.538 0.458
T-NeRF+B+D [15] 17.06 0.573 0.390
T-NeRF+B+D+S [15] 16.96 0.577 0.379
RoDynRF [27] 17.00 0.535 0.517
4D-GS [55] 13.87 0.459 0.407
Ours 15.79 0.585 0.354
[…]
We compare the performance of several related methods in Tab. 1. For all NeRF-based approaches, we report values from the literature. To further compare against a fast forward rendering approach, we train the state-of-the-art Gaussian splatting-based 4D-GS [55] on both benchmark datasets. For both 4D-GS and our method, we measure performance on a single RTX 4090 GPU. As evident from the results, our method is significantly faster than backward rendering approaches, and even trains faster than 4D-GS while achieving similar frame rates during inference. A breakdown of timings for individual method components is provided in the supplemental material. We further compare image quality based on PSNR, SSIM [54] and LPIPS [62] quality metrics on the NVIDIA and iPhone datasets in Tab. 2. To avoid inconsistencies across the reported values, we use the official results and models kindly provided by Liu et al. [27], Li et al. [26], and Gao et al. [15] to consistently recalculate all metrics using the VGG backbone for LPIPS calculation. For the iPhone dataset, we use masked metric calculation (mPSNR, mSSIM, mLPIPS) as suggested and implemented by Gao et al. [15] to account for co-visibility. Overall, we find our method is not only the fastest in terms of reconstruction and rendering speed but also achieves competitive image quality, especially in terms of perceptual image quality metrics (LPIPS). While RoDynRF and T-NeRF achieve better per-pixel PSNR scores, we find that both methods fail to preserve high frequencies, resulting in an inferior perceived image quality, as shown in the next section. Detailed per-scene results can be found in our supplement.
4.2 Qualitative Results
TiNeuVox RoDynRF 4D-GS Ours Ground Truth
Figure 2: Qualitative comparisons on the NVIDIA dataset. Our method preserves fine details and clean foreground-background transitions.
[…]
Figure 3: Comparison on the iPhone dataset. We compare on two challenging scenes of the iPhone dataset. Our method better preserves the object shapes and fine details despite complex deformations. Methods indicated by $\dagger$ are the highest quality versions provided by Gao et al. [15]
We show qualitative results on the NVIDIA (Fig. 2) and iPhone (Fig. 3) datasets. For comparisons on the iPhone sequences, we show the best version of HyperNeRF [35] and T-NeRF [37] according to image quality metrics, using all additional regularizers (Lidar depth, random background, and distortion loss) proposed by Gao et al. [15]. For our method, we report values for two versions, trained with the provided Lidar or monocular depth data. Again, we find that our method achieves a competitive and oftentimes superior perceived image quality, preserving higher frequencies and clean edges. For dynamic view synthesis results, please see our supplemental video.
4.3 Ablation Study
Table 3: Comparison of ablated D-NPC versions in terms of image quality, model size and speed.
Method Image Quality Model Size Frame Rate
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ [MB] $\downarrow$ [FPS] $\uparrow$
Oursfull 25.64 0.845 0.109 668 72
Oursw/o init 21.70 0.617 0.302 813 45
Oursw/o depth 24.40 0.835 0.135 671 72
Oursw/o U-Net 24.24 0.798 0.155 662 78
To assess the improvements enabled by our main pipeline components, we optimize three variations of our method on the NVIDIA dataset, and compare them to our full D-NPC model (full). First, we train our model from scratch without initializing the probability field from the monocular depth and foreground segmentation, just using camera frustum carving to reduce the amount of initial cells (w/o init). Additionally, we compare against a version without depth supervision during training (w/o depth), which is a crucial component in monocular dynamic view synthesis. Lastly, we evaluate the impact of our neural renderer by skipping the U-Net, directly interpreting the rasterized point cloud features as RGB outputs (w/o U-Net). As evident from Tab. 3, our initialization procedure effectively guides the model to a favorable representation, decreasing image quality and increasing memory when omitted. Likewise, skipping additional depth supervision from the globally aligned monocular depth maps significantly deteriorates reconstruction quality. The neural rendering network used for hole filling, while also contributing to improved final image quality, comes at the cost of a slightly decreased inference frame rate. We find that direct RGB rasterization without a rendering network can still produce visually appealing results, and improves the interactivity at higher resolutions when exploring scenes in a graphical user interface. Importantly, the U-Net weights do not have any significant impact on the final model size, which is dominated by the dynamic sampling and appearance parameters.
5 Discussion and Limitations
Figure 4: Failure Cases. Top: Cropped foreground not contained in the training camera frustum. Bottom: Object duplications due to depth misalignment.
In our experiments, our D-NPC yields the best trade-off between optimization time and image quality and is currently the only monocular method providing real-time frame rates for interactive applications. Especially in terms of perceptual LPIPS and high image frequencies, our method quantitatively and visually outperforms recent monocular view synthesis methods while only requiring a fraction of the optimization time. However, there is room for improvement: Our method relies on a decent quality of camera poses and monocular depth. Moreover, our initialization process assumes that the dynamic content is fully contained in the camera frustum, resulting in visible crops when synthesizing far-off novel views, as shown in Fig. 4.
6 Conclusion
We presented Dynamic Neural Point Clouds, an efficient implicit point cloud-driven [17] approach for non-rigid novel-view synthesis from monocular video inputs. In contrast to existing ray marching-based NeRF methods, our rasterization-based forward rendering approach enables a direct and elegant way to initialize and guide the reconstruction from data-driven priors, not only making it one of the fastest approaches in terms of optimization time, but also enabling rendering at real-time frame rates. While our experiments show our novel approach already yields competitive image quality according to several metrics on common benchmark datasets, we believe that our work constitutes a promising new direction for future research and extensions, like the integration of diffusion-based priors.
Acknowledgments and Disclosure of Funding
We thank Jann-Ole Henningson for help with figures and writing.
This work was partially funded by the German Research Foundation (DFG) projects “Immersive Digital Reality” (ID 283369270) and “Real-Action VR” (ID 523421583).
Appendix A Appendix – Supplemental Material
In this supplemental, we provide details on our model configuration (Sec. A.1) and optimization procedure (Sec. A.2), include additional results (Sec. A.3), and discuss concurrent work on arXiv (Sec. A.4) as well as the societal impact (Sec. A.5) of our work.
A.1 Model Details
Table 4: Appearance Field Hash Grid Configurations.
Parameter 3D Static Grid 4D Dynamic Grid
Base Resolution 16 16
Resolution Levels 10 8
Per Level Scale 2 2
Features Per Level 4 4
Logarithmic Size 21 22
In our model, the crucial neural components regarding the image quality are the appearance feature grids and the neural rendering network. To model our static and dynamic appearance fields, we use the fast multi-resolution hash encoding implementation by Müller et al. [33, 32] After initially carving our sampling distribution voxel grid using the monocular depth and dynamic segmentation priors, we set the spatial scale of our 3D and 4D feature grids to the minimum axis-aligned bounding box enclosing all remaining static and dynamic voxels, respectively. We parameterize the hash-grids according to the values in Tab. 4, giving more total capacity to the higher dimensional dynamic representation. Both hash encodings are followed by a shallow ReLU-activated MLP with one hidden $64$-neuron layer to resolve hash collisions.
For our neural rendering network used for hole-filling, we apply a simple image-to-image translation approach based on a U-Net architecture [38]. In contrast to Hahlbohm et al. [17], we do not use any multi-resolution inputs. Instead, we use two downsampling and consecutive upsampling blocks, each consisting of two ReLU-activated 2D convolutions with kernel size $3$, followed by average pooling / bilinear upsampling. During down- and upscaling, each convolution block doubles/halves the amount of feature channels while adjusting the spatial resolution, going from an initial $64$ channels at the finest to a maximum of $256$ channels at the coarsest spatial resolution.
A.2 Optimization Details
Table 5: Loss Hyperparameters.
Hyperparameter Value
$\lambda_{\text{photo, Cauchy}}$ 1.0
$\lambda_{\text{photo, D-SSIM}}$ 0.9
$\lambda_{\text{photo, LPIPS}}$ 0.05
$\lambda_{\text{depth}}$ 20.0
$\lambda_{\text{seg}}$ $1\mathrm{e}{-3}$
$\lambda_{\text{dist}}$ 1.0
$k$ 0.5
We train our model for a total of 10K iterations on a consumer-grade RTX 4090 GPU. During each iteration, we sample an explicit point cloud consisting of $4$M points (i.e., a batch of spatial 3D locations). While the sampling distribution is manually updated based on statistics of the extracted explicit point clouds (see Sec. 3.2.1), the differentiable rasterizer and neural rendering network enable end-to-end optimization using gradient descent for the network and hash-grid parameters. We use the Adam [20] optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1\mathrm{e}{-15}$) with an initial learning rate of $1\mathrm{e}{-2}$ for the appearance hash-grids and $5\mathrm{e}{-4}$ for the rendering network, which are exponentially decayed to $\frac{1}{3}\mathrm{e}{-3}$ and $5\mathrm{e}{-5}$ during optimization. For our loss hyperparameters (Sec. 3.4), we use the values provided in Tab. 5. Note that, after initially training with a high value $\lambda_{depth}$ to guide the optimization in a reasonable direction, we turn off our depth loss after $500$ training iterations to allow the model to recover from errors in the monocular depth estimates. In contrast to D2NeRF, who use $k=2$ to bias the binary segmentation loss towards static background, which usually constitutes a majority of the scene, we use a value of $0.5$ to favor the dynamic model. This is due to the fact that our initialization procedure already discards large amounts of dynamic model evaluation in static areas. As a result, the remaining dynamic samples are already biased towards potentially dynamic areas, which we reflect in the choice of skew $k$. In addition to the losses already discussed in Sec. 3.4, we further apply weight decay (i.e., an L2 loss) with a factor of $2.5$ on the appearance grid parameters [5]. In practice, we find that this decay helps to prevent overfitting and improves the quality of generated novel views. Our Python / CUDA implementation will be made publicly available.
A.3 Additional Results
We provide detailed breakdowns on the NVIDIA and iPhone datasets in Tab. 6 and Tab. 7, respectively. As evident from the per-scene results, our method consistently achieves competitive results across a variety of challenging scenes, and particularly excels at perceptual LPIPS scores due to its ability to retain high-frequency details.
Table 6: Quantitative comparison on the NVIDIA dataset. We compare our method against recent dynamic view synthesis approaches on the NVIDIA dynamic scenes dataset. Our method achieves the best scores in terms of perceptual image quality. The three best results are highlighted in green in descending order of saturation.
Balloon1 Balloon2 Jumping Playground
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
T-NeRF [37] 18.54 0.448 0.416 20.69 0.573 0.335 18.04 0.496 0.541 14.68 0.240 0.527
D-NeRF [37] 19.06 0.492 0.382 20.76 0.557 0.367 22.36 0.712 0.318 20.18 0.670 0.250
HyperNeRF [35] 13.96 0.214 0.677 16.57 0.286 0.544 18.34 0.521 0.460 13.17 0.179 0.603
NR-NeRF [49] 17.39 0.369 0.497 22.41 0.703 0.314 20.09 0.644 0.418 15.06 0.248 0.450
NSFF [25] 21.96 0.701 0.281 24.27 0.731 0.276 24.65 0.813 0.227 21.22 0.705 0.263
DynamicNeRF [14] 22.36 0.775 0.169 27.06 0.859 0.113 24.68 0.842 0.153 24.15 0.849 0.145
RoDynRF [27] 22.37 0.782 0.157 26.19 0.846 0.107 25.66 0.853 0.129 24.96 0.899 0.080
TiNeuVox [11] 17.30 0.392 0.500 19.06 0.458 0.396 20.81 0.641 0.399 13.84 0.213 0.541
4D-GS [55] 20.82 0.666 0.271 24.34 0.775 0.211 22.16 0.735 0.270 20.22 0.724 0.192
Ours 22.26 0.779 0.161 25.84 0.835 0.107 24.51 0.839 0.116 23.59 0.866 0.099
Skating Truck Umbrella Average
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
T-NeRF [37] 20.32 0.551 0.562 18.33 0.462 0.523 17.69 0.284 0.670 18.33 0.436 0.511
D-NeRF [37] 22.48 0.678 0.403 24.10 0.743 0.213 21.47 0.554 0.310 21.49 0.629 0.320
HyperNeRF [35] 21.97 0.640 0.322 20.61 0.493 0.394 18.59 0.308 0.517 17.60 0.377 0.502
NR-NeRF [49] 23.95 0.745 0.342 19.33 0.485 0.548 19.63 0.376 0.471 19.69 0.510 0.434
NSFF [25] 29.29 0.889 0.186 25.96 0.775 0.248 22.97 0.644 0.315 24.33 0.751 0.257
DynamicNeRF [14] 32.66 0.951 0.077 28.56 0.872 0.146 23.26 0.711 0.190 26.10 0.837 0.142
RoDynRF [27] 28.68 0.939 0.076 29.13 0.900 0.089 24.26 0.757 0.135 25.89 0.854 0.110
TiNeuVox [11] 23.32 0.723 0.283 23.86 0.662 0.328 20.00 0.416 0.439 19.74 0.501 0.412
4D-GS [55] 26.04 0.876 0.171 25.08 0.753 0.233 21.59 0.588 0.309 22.89 0.731 0.237
Ours 30.22 0.948 0.061 28.92 0.897 0.084 24.15 0.747 0.136 25.64 0.845 0.109
Table 7: Quantitative evaluation on the iPhone dataset. We compare our method against related approaches using masked metrics based on co-visibility [15]. Methods marked with $\dagger$ are trained using the code and additional regularizers introduced by Gao et al. [15]. We further show values for our method trained on monocular depth instead of the provided Lidar data. The three best results are highlighted in green in descending order of saturation.
Apple Block Paper-windmill Space-out
Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
HyperNeRF+B+D+S [35] 17.64 0.743 0.478 17.54 0.670 0.331 17.38 0.382 0.209 17.93 0.605 0.320
Nerfies+B+D+S [34] 17.54 0.750 0.478 16.61 0.639 0.389 17.34 0.378 0.211 17.79 0.622 0.303
T-NeRF [15] 16.64 0.719 0.559 16.36 0.637 0.434 15.29 0.278 0.447 17.17 0.602 0.397
T-NeRF+B+D [15] 17.32 0.720 0.514 17.60 0.670 0.375 17.56 0.351 0.295 18.34 0.619 0.345
T-NeRF+B+D+S [15] 17.43 0.728 0.508 17.52 0.669 0.346 17.55 0.367 0.258 17.71 0.591 0.377
RoDynRF [27] 18.74 0.723 0.552 18.05 0.634 0.513 16.71 0.321 0.482 18.57 0.594 0.413
4D-GS [55] 15.35 0.686 0.432 13.40 0.544 0.506 14.48 0.201 0.336 14.37 0.513 0.366
Ours 16.45 0.749 0.487 15.32 0.626 0.363 18.03 0.434 0.209 17.17 0.624 0.288
[…]
Spin Teddy Wheel Average
Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
HyperNeRF+B+D+S [35] 19.20 0.561 0.325 13.97 0.568 0.350 13.99 0.455 0.310 16.81 0.569 0.332
Nerfies+B+D+S [34] 18.38 0.585 0.309 13.65 0.557 0.372 13.82 0.458 0.310 16.45 0.570 0.339
T-NeRF [15] 16.36 0.504 0.530 13.38 0.550 0.475 14.73 0.477 0.362 15.70 0.538 0.458
T-NeRF+B+D [15] 18.95 0.541 0.468 13.99 0.573 0.435 15.69 0.541 0.301 17.06 0.573 0.390
T-NeRF+B+D+S [15] 19.16 0.567 0.443 13.71 0.570 0.429 15.65 0.548 0.292 16.96 0.577 0.379
RoDynRF [27] 17.41 0.484 0.570 14.34 0.537 0.613 15.20 0.449 0.478 17.00 0.535 0.517
4D-GS [55] 15.36 0.418 0.310 12.36 0.506 0.474 11.78 0.343 0.427 13.87 0.459 0.407
Ours 17.07 0.567 0.322 12.63 0.541 0.475 13.85 0.555 0.330 15.79 0.585 0.354
[…]
Table 8: Foreground analysis on the short NVIDIA dataset. We show additional quantitative analysis on the short NVIDIA sequences. We use masked metrics [15] with the provided foreground masks to evaluate image quality on only the dynamic part of the scenes. The three best results are highlighted in green in descending order of saturation.
Balloon1 Balloon2 Jumping Playground
Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
T-NeRF [37] 17.90 0.897 0.245 19.89 0.953 0.133 15.24 0.914 0.285 14.81 0.957 0.267
D-NeRF [37] 18.75 0.907 0.294 18.23 0.950 0.203 19.23 0.941 0.160 16.65 0.970 0.227
HyperNeRF [35] 14.43 0.865 0.471 15.95 0.931 0.251 14.82 0.913 0.278 12.32 0.950 0.390
NR-NeRF [49] 15.87 0.880 0.383 20.13 0.962 0.177 16.76 0.934 0.228 13.91 0.956 0.311
NSFF [25] 19.34 0.925 0.225 22.63 0.969 0.155 20.42 0.955 0.152 20.22 0.983 0.162
DynamicNeRF [14] 19.73 0.931 0.134 23.79 0.981 0.042 20.68 0.958 0.100 21.05 0.984 0.095
RoDynRF [27] 19.70 0.928 0.147 23.05 0.977 0.046 21.54 0.962 0.093 20.29 0.984 0.097
TiNeuVox [11] 16.80 0.888 0.330 19.59 0.955 0.147 17.34 0.932 0.196 13.02 0.950 0.363
4D-GS [55] 18.72 0.911 0.247 22.21 0.973 0.066 18.60 0.938 0.173 17.23 0.976 0.156
Ours 19.76 0.930 0.159 22.14 0.973 0.056 21.07 0.960 0.080 19.76 0.982 0.107
Skating Truck Umbrella Average
Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
T-NeRF [37] 22.56 0.993 0.086 18.40 0.913 0.409 16.32 0.945 0.496 17.87 0.939 0.274
D-NeRF [37] 22.05 0.993 0.090 19.84 0.923 0.312 17.34 0.953 0.365 18.87 0.948 0.236
HyperNeRF [35] 23.86 0.993 0.074 19.40 0.918 0.281 14.76 0.940 0.489 16.51 0.930 0.319
NR-NeRF [49] 27.17 0.995 0.052 18.25 0.912 0.534 16.23 0.948 0.404 18.33 0.941 0.298
NSFF [25] 31.37 0.998 0.019 26.51 0.971 0.135 19.28 0.967 0.238 22.82 0.967 0.155
DynamicNeRF [14] 37.22 0.999 0.010 26.75 0.973 0.067 19.66 0.969 0.152 24.13 0.971 0.086
RoDynRF [27] 34.37 0.997 0.035 27.29 0.978 0.061 20.85 0.970 0.134 23.87 0.971 0.087
TiNeuVox [11] 27.18 0.996 0.047 21.39 0.934 0.204 16.09 0.949 0.358 18.77 0.943 0.235
4D-GS [55] 30.99 0.997 0.048 22.77 0.947 0.207 18.36 0.956 0.259 21.27 0.957 0.165
Ours 36.38 0.999 0.014 27.16 0.975 0.063 21.71 0.975 0.087 24.00 0.971 0.081
As most natural scenes are dominated by static background, we individually assess the quality of dynamic foreground reconstruction, which is the most complex part of dynamic view synthesis from monocular inputs. To this end, we recalculate metrics on the NVIDIA sequences using masked metrics [15], based on the motion masks contained in the dataset. As shown in Tab. 8, our method still performs favorably regarding perceptual metrics, indicating that improvements were not exclusively located in the static background.
As our method can handle significantly longer and temporally denser sequences than the commonly used train/test splits of the NVIDIA dynamic scenes dataset, we further compare against DynIBaR [26], using their evaluation protocol containing $\sim 150$ training views per scene. For fairness, we also use the consistent video depth made available by the authors. Results are shown in  Tab. 9.
Table 9: Comparison on the long NVIDIA dataset. We compare to the state-of-the-art approach DynIBaR [26] on the long sequences from the NVIDIA dynamic scenes dataset. Masked metrics are calculated with respect to the dynamic foreground masks.
Method Balloon1 Balloon2 Jumping Playground Skating Truck Umbrella DynamicFace Average
PSNR $\uparrow$
DynIBaR [26] 28.99 28.90 23.74 28.32 31.50 33.75 28.25 29.18 29.08
Ours 27.32 30.09 22.67 25.51 33.29 32.14 24.76 28.18 28.00
SSIM $\uparrow$
DynIBaR [26] 0.920 0.942 0.866 0.953 0.969 0.966 0.898 0.971 0.936
Ours 0.902 0.934 0.841 0.895 0.969 0.952 0.737 0.962 0.899
LPIPS $\downarrow$
DynIBaR [26] 0.064 0.070 0.126 0.049 0.058 0.049 0.085 0.052 0.069
Ours 0.065 0.052 0.128 0.072 0.044 0.045 0.142 0.048 0.074
mPSNR $\uparrow$
DynIBaR [26] 23.91 25.53 17.49 20.91 19.52 29.06 25.96 30.55 24.12
Ours 22.35 23.51 16.30 18.30 19.89 25.57 21.75 23.94 21.45
mSSIM $\uparrow$
DynIBaR [26] 0.944 0.985 0.911 0.989 0.990 0.988 0.971 0.997 0.972
Ours 0.937 0.980 0.896 0.984 0.992 0.979 0.921 0.991 0.960
mLPIPS $\downarrow$
DynIBaR [26] 0.079 0.032 0.168 0.062 0.070 0.038 0.043 0.010 0.062
Ours 0.077 0.037 0.193 0.097 0.069 0.053 0.110 0.026 0.083
While our method does not quite attain the image quality of DynIBaR, we achieve competitive results with significantly lower training and rendering times.
Table 10: Timings for individual method components.
      Component       Time [s]
      Data Preprocessing       140
      Depth Alignment       6
      Model Initialization       1
      Model Optimization       817
We further provide a breakdown of timings for individual components of our method on the NVIDIA dataset in Tab. 10. Note that the cost of data preprocessing highly depends on the amount of input images, while the optimization and rendering times increase with image resolution.
A.4 Concurrent Work
Concurrent to our work, other methods on arXiv address dynamic reconstruction and novel-view synthesis from monocular video. Lee et al. [23] propose a new scene representation for fast reconstruction and interactive rendering from in-the-wild monocular videos. Similar to our approach, they separately model the foreground and background, but use plane based proxy geometry for static regions, and point clouds for dynamic foreground, enabling fast forward rendering and an impressive image quality. More details and results are presented at the official project page: https://casual-fvs.github.io.
Following recent trends in static scene reconstruction, Wang et al. [50] integrate diffusion-based priors in the monocular reconstruction. While not being exceptionally fast, their approach hallucinates additional training supervision for occluded areas, improving results on the challenging iPhone dataset [15]. Again, we refer the interested reader to the official project page for more information: https://mightychaos.github.io/dpdy_proj/.
A.5 Societal Impact
Our method reconstructs a 4D representation and generates novel views based on a monocular input video. As the method itself cannot fill unseen areas or use/provide any information that goes beyond the inputs provided by the users themselves, we do not expect any negative societal impact resulting from our work. In terms of environmental impact, the short training times and consumer-grade GPU (RTX 3090/4090) required by our method makes it very economic in terms of power consumption and hardware requirements: Overall, after downloading and setting up our code, running a single reconstruction, including data preprocessing with user inputs, takes under one hour of time. When running on a slightly above average PSU with $750$W at full capacity (which is usually not the case), given an average energy price of \$$0.173$ per kWh (April 2024 in the U.S.), we get a worst-case cost of \$$0.13$ per scene reconstruction. Thus, in comparison to DynIBaR [26], the current the state-of-the-art in terms of image quality, which requires $384$ GPU hours on high-end A100 GPUs, our method has a significantly lower environmental footprint.

### v2

@@
-D-NPC: Dynamic Neural Point Clouds for Non-Rigid View Synthesis from Monocular Video
+
 Abstract
@@
 As evidenced by our experimental evaluation, our dynamic point cloud model not only enables fast optimization and real-time frame rates for interactive applications, but also achieves competitive image quality on monocular benchmark sequences.
-Our project page is available at https://moritzkappel.github.io/projects/dnpc.
+Our code and data are available online https://moritzkappel.github.io/projects/dnpc/.
+{CCSXML}
+<ccs2012> <concept> <concept_id>10010147.10010371.10010382.10010385</concept_id> <concept_desc>Computing methodologies Image-based rendering</concept_desc> <concept_significance>100</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010396.10010400</concept_id> <concept_desc>Computing methodologies Point-based models</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010178.10010224.10010245.10010254</concept_id> <concept_desc>Computing methodologies Reconstruction</concept_desc> <concept_significance>100</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010372.10010373</concept_id> <concept_desc>Computing methodologies Rasterization</concept_desc> <concept_significance>300</concept_significance> </concept> </ccs2012>
 […]
-figure We introduce Dynamic Neural Point Clouds for dynamic view synthesis from monocular videos. Our method enables fast optimization, real-time frame rates, and competitive image quality.
+[…]
+[…]
 1 Introduction
@@
 Using recent advances in 3D neural scene reconstruction, it becomes possible to create photorealistic renderings from arbitrary viewpoints, effectively adding interactive six degrees-of-freedom camera support.
-In the context of general, non-rigidly deforming dynamic scenes, the capability of interactively rendering novel views for every point in time enables numerous applications, including video stabilization [28], immersive playback for social media [2], or virtual reality [57].
+In the context of general, non-rigidly deforming dynamic scenes, the capability of interactively rendering novel views for every point in time enables numerous applications, including video stabilization [LLY21], immersive playback for social media [AHR23], or virtual reality [XAL23].
 However, performing faithful reconstruction from just a single monocular video (e.g., a casual smartphone recording) remains a computationally expensive and challenging problem.
-While the introduction of Neural Radiance Fields [31] and 3D Gaussian Splatting [19] brought about significant advances in static scene reconstruction, dynamic extensions still lag behind in terms of quality and robustness due to motion and depth ambiguities in the temporal domain.
-Those ambiguities are usually resolved using expensive large-scale multi view setups [24] or monocularized data (i.e., teleporting camera setups) [18], which maximize the ratio between camera and scene motion to improve temporal consistency.
-However, recent investigations [15] reveal that most sparsely regularized methods fail to faithfully recover consistent motion and geometry from casual smartphone captures due to the lack of effective multi-view signal.
-Multiple recent radiance field approaches address this problem by incorporating strong priors like monocular depth, optical flow, and semantic segmentation for the joint reconstruction of geometry, 4D scene flow fields [25, 26], and even camera parameters [27] from monocular video.
-While current monocular methods achieve high-quality spatiotemporally consistent view synthesis, they are often limited in the expressiveness of novel camera views, and by their high computational demand.
+While the introduction of Neural Radiance Fields [MST21] and 3D Gaussian Splatting [KKLD23] brought about significant advances in static scene reconstruction, dynamic extensions still lag behind in terms of quality and robustness due to motion and depth ambiguities in the temporal domain.
+Those ambiguities are usually resolved using expensive large-scale multi view setups [LSZ22] or monocularized data (i.e., teleporting camera setups) [KGC25], which maximize the ratio between camera and scene motion to improve temporal consistency.
+However, recent investigations [GLT22] reveal that most sparsely regularized methods fail to faithfully recover consistent motion and geometry from casual smartphone captures due to the lack of effective multi-view signal.
+Multiple recent radiance field approaches address this problem by incorporating strong priors like depth, optical flow, and semantic segmentation for the joint reconstruction of geometry, 4D scene flow fields [LNSW21, LWC23], and even camera parameters [LGM23] from monocular video.
+While current monocular methods achieve high-quality spatio-temporally consistent view synthesis, they are often limited in the expressiveness of novel camera views, and by their high computational demand.
 To address the above-mentioned challenges, this paper introduces Dynamic Neural Point Clouds (D-NPC), a new system for monocular dynamic view synthesis based on implicit point distributions and fast hash-encoded feature grids.
 Given a monocular input video and the corresponding structure-from-motion calibration, our method optimizes a 4D representation of the depicted scene, which enables spatio-temporal novel-view synthesis, e.g., the infamous bullet-time effect.
-Following the concept of implicit neural point clouds [17], we track expected scene density in a sparse voxel grid, which is used to extract explicit pose-dependent point clouds.
-Combined with a fast multi-resolution hash-encoded feature grid [33], this scene representation is able to efficiently render high-fidelity novel views using a differentiable rasterizer and image-to-image translation network.
-Based on the concepts of space-time hash encoding [51], we lift the representation to the temporal domain by replacing the single feature grid with a 3D and 4D hash encoding for the static background and dynamic foreground regions, respectively.
+Following the concept of implicit neural point clouds [HFK24], we track expected scene density in a sparse voxel grid, which is used to extract explicit pose-dependent point clouds.
+Combined with a fast multi-resolution hash-encoded feature grid [MESK22], this scene representation is able to efficiently render high-fidelity novel views using a differentiable rasterizer and image-to-image translation network.
+Based on the concepts of space-time hash encoding [WCW23], we lift the representation to the temporal domain by replacing the single feature grid with a 3D and 4D hash encoding for the static background and dynamic foreground regions, respectively.
 We further extend the point distribution field to robustly track the per-voxel ratio of static and dynamic scene content, effectively enabling 3D dynamics segmentation to significantly reduce the required compute during optimization and inference.
-To guide the severely under-constrained monocular reconstruction, we leverage per-frame data-driven priors, namely optical flow, monocular depth, and foreground segmentation masks, which we extract and preprocess with minimal overhead.
+To guide the severely under-constrained monocular reconstruction, we leverage per-frame data-driven priors, namely monocular depth and foreground segmentation masks, which we extract and preprocess with minimal overhead.
 We demonstrate that our temporal extension can take advantage of the rasterization-based forward rendering of implicit neural point clouds, which enables efficient foreground-background decomposition and image-space regularization.
@@
 Rendering novel viewpoints of a recorded scene is a longstanding and well-studied problem.
-Early work on image-based rendering uses morphing to generate in-between views from image collections without explicit 3D reconstruction [9, 42].
-Other approaches apply implicit or explicit 3D scene representations like light fields [16] or layered depth images [43] to generate high-quality novel views.
-Recently, Neural Radiance Fields (NeRFs) [31] and 3D Gaussian Splatting (3D-GS) [19] inspired a multitude of follow-up research, including extensions for dynamic scenes.
+Early work on image-based rendering uses morphing to generate in-between views from image collections without explicit 3D reconstruction [CW93, SD96].
+Other approaches apply implicit or explicit 3D scene representations like light fields [GGSC96], per-view meshes [HPP18], geometric scaffolds [RK21], or layered depth images [SGHS98] to generate high-quality novel views.
+Recently, Neural Radiance Fields (NeRFs) [MST21] and 3D Gaussian Splatting (3D-GS) [KKLD23] inspired a multitude of follow-up research, including extensions for dynamic scenes.
 We next briefly summarize neural representations for static scenes, followed by dynamic view synthesis.
 2.1 Static View Synthesis
-NeRFs [31] implicitly represent the scene geometry and appearance using a Multi-Layer Perceptron (MLP), yielding an unprecedented image quality for novel views at the cost of slow optimization and rendering due to the large number of network queries required.
-While some follow-ups propose qualitative improvements for, e.g., anti-aliasing and unbounded scenes [3, 4, 5], another branch of methods focuses on acceleration using explicit scene representations like sparse [13] or decomposed [8] voxel grids.
+NeRFs [MST21] implicitly represent the scene geometry and appearance using a Multi-Layer Perceptron (MLP), yielding an unprecedented image quality for novel views at the cost of slow optimization and rendering due to the large number of network queries required.
+While some follow-ups propose qualitative improvements for, e.g., anti-aliasing and unbounded scenes [BMT21, BMV22, BMV23], another branch of methods focuses on acceleration using explicit scene representations like sparse [FKYT22] or decomposed [CXG22] voxel grids.
 Along the same lines, Müller et al.
-[33] propose an efficient grid-based scene representation implemented as a multi-resolution hash-grid, enabling high-quality reconstruction in only a few minutes.
+[MESK22] propose an efficient grid-based scene representation implemented as a multi-resolution hash-grid, enabling high-quality reconstruction in only a few minutes.
 All the above volume rendering approaches apply backward rendering, where the scene is frequently sampled along the camera rays.
@@
 The popular 3D-GS by Kerbl et al.
-[19] represents the scene as a set of 3D Gaussians for very fast rendering at a high image quality, inspiring many extensions [22, 60].
-Analogously, point-based rendering approaches [1, 21, 39] use a fixed initial point cloud combined with a neural image translation network for hole filling.
+[KKLD23] represents the scene as a set of 3D Gaussians for very fast rendering at high image quality, inspiring many extensions [LRS24, YCH24].
+Analogously, point-based rendering approaches [ASK20, KPLD21, RFS22] use a fixed initial point cloud combined with a neural image translation network for hole filling.
 The recent INPC approach by Hahlbohm et al.
-[17] combines the benefits of forward and backward rendering, modelling the scene as a point probability and hash-encoded appearance field for high-quality rendering at interactive frame rates.
+[HFK24] combines the benefits of forward and backward rendering, modelling the scene as a point probability and hash-encoded appearance field for high-quality rendering at interactive frame rates.
 Our D-NPC method leverages these benefits for fast dynamic view synthesis from monocular video.
 A more detailed discussion of neural rendering techniques can be found in the report by Tewari et al.
-[48].
+[TTM22].
 2.2 Dynamic View Synthesis
-Most NeRF-based extensions use a time-conditioned MLP [24, 25, 36] or additional deformation field [11, 34, 35, 37] to represent scene dynamics.
-Following acceleration models for static reconstruction, similar models based on factorized [2, 7, 12, 52] and hash-encoded [18, 45, 51] feature grids were introduced for dynamic view synthesis.
-D2NeRF [56] combined two models for foreground-background separation from a single video.
+Most NeRF-based extensions use a time-conditioned MLP [LSZ22, LNSW21, PSJ23] or additional deformation field [FYW22, PSB21, PSH21, PCPMMN21] to represent scene dynamics.
+Following acceleration models for static reconstruction, similar models based on factorized [AHR23, CJ23, FKMW23, WTL23] and hash-encoded [KGC25, SCL23, WCW23] feature grids were introduced for dynamic view synthesis.
+D2NeRF [WZT22] combined two models for foreground-background separation from a single video.
 Likewise, Wang et al.
-[51] decompose the scene in a fast 3D and 4D hash-grid representation.
+[WCW23] decompose the scene in a fast 3D and 4D hash-grid representation.
 We adapt concepts from both approaches, integrating foreground/background handling in both our sampling and appearance modelling, enabling our method to individually render dynamic and static scene content while improving performance.
-Similar to radiance fields, recent approaches opt to extend 3D-GS for non-rigid scenes, tracking the 3D motion of Gaussians over time [29, 55] while retaining fast rendering.
+Similar to radiance fields, recent approaches opt to extend 3D-GS for non-rigid scenes, tracking the 3D motion of Gaussians over time [LKLR24, WYF24] while retaining fast rendering.
 However, as recently revealed by Gao et al.
-[15], these methods require multi-view or monocularized [18] capturing setups to exploit observations from multiple views.
-On the other hand, multiple methods focus on dynamic view synthesis from a single monocular input video, integrating additional priors like dense scene flow estimation to resolve motion and depth ambiguities [25, 26, 27].
-RoDynRF [27] further takes the challenging scenario to the extreme, additionally optimizing camera poses from the unposed input video.
+[GLT22], these methods require multi-view or monocularized [KGC25] capturing setups to exploit observations from multiple views.
+On the other hand, multiple methods focus on dynamic view synthesis from a single monocular input video, integrating additional priors like dense scene flow estimation to resolve motion and depth ambiguities [LNSW21, LWC23, LGM23].
+RoDynRF [LGM23] further takes the challenging scenario to the extreme, additionally optimizing camera poses from the unposed input video.
 However, these MLP-based monocular approaches rely on backward rendering, thus being limited in optimization speed and final rendering frame rate.
-This paper presents a point-based forward rendering approach for monocular view synthesis, which leverages additional priors to initialize and guide the reconstruction, enabling high-quality rendering at real-time frame rates.
+To circumvent slow optimization, the PGDVS approach by Zhao et al.
+[ZCM24] proposes a generalized approach for monocular video that bypasses per scene optimization.
+Still, the slow inference speed and varying generalization quality largely rule out application for real-time downstream tasks.
+To jointly enable fast high quality per-scene optimization and subsequent real-time novel-view synthesis, this paper presents a point-based forward rendering approach for monocular view synthesis, which leverages additional priors to initialize and guide the reconstruction.
 More details on the state of the art in dynamic reconstruction can be found in the survey by Yunus et al.
-[61].
+[YLN24].
+2.3 Concurrent Work
+Concurrent to our work, other approaches propose fast forward rendering representations for dynamic reconstruction and novel-view synthesis from monocular video.
+Please note that the covered concurrent methods are not yet published in any proceedings (at the time of submission), or constitute non-peer reviewed work on arXiv.
+Similar to our approach, Lee et al.
+[LZBM25] separately model the foreground and background, but instead use a plane based proxy geometry for static regions, and points for the dynamic foreground, enabling fast forward rendering at a decent image quality.
+Following up on the adaption of 3DGS for dynamic multi-view recordings, Stearns et al.
+[SHU24] propose a divide-and-conquer approach for motion and appearance reconstruction from single monocular video using isotropic Gaussians.
+Like most Gaussian-based methods, this approach inherits impressive online framerates and decent image quality, but adds overhead in terms of optimization time.
+The Shape of Motion approach by Wang et al.
+[WYG24] additionally uses 2D motion track inputs to recover 3D scene motion represented via decomposed motion bases, which is applied to a set of canonical 3D Gaussians for novel view synthesis.
+Lei et al.
+[LWH24] also propose to disentangle the deformations of the scene from its geometry.
+The former is represented as a sparse graph of motion trajectories that they lift from 2D priors and optimize exploiting semantic features and the assumption of the deformations being as-rigid-as-possible.
+They use Gaussian Splatting to render the scene resulting from the global fusion of the 3DGs per timestep and deformed accordingly to the motion graph.
+While our method also involves rendering explicit geometry for improved performance, using an implicit point cloud representation instead of explicit Planes, Points, or Gaussians entails unique characteristics regarding performance, quality and robustness.
 3 Method
@@
 We introduce Dynamic Neural Point Clouds (D-NPC), a method for dynamic novel-view synthesis from monocular video.
-Given a single monocular video consisting of $N$ images $I=(i_{1},...,i_{N})$ and their normalized timestamps $T=(t_{1},...,t_{N})\in[0,1]^{N}$, we reconstruct a dynamic implicit point cloud representation [17] of the depicted dynamic scene.
+Given a single monocular video consisting of $N$ images $I=(i_{1},...,i_{N})$ and their normalized timestamps $T=(t_{1},...,t_{N})\in[0,1]^{N}$, we reconstruct a dynamic implicit point cloud representation [HFK24] of the depicted dynamic scene.
 Based on its static counterpart, the core of our D-NPC model consists of a spatiotemporal point probability ($\mathcal{P}$) and appearance feature descriptor ($\mathcal{A}$) field, which can efficiently be rendered using a differentiable rasterizer and neural rendering network.
@@
 The reconstruction of non-rigidly deforming scenes from a single moving camera is a highly ill-posed problem.
-To regularize the reconstruction from monocular video, previous approaches [15, 25, 27] apply estimates from several image processing applications, replacing the multi-view consistency signal with learned, data-driven scene analysis.
-Following this idea, our method starts with data preprocessing to extract high-quality priors, which we use to initialize and guide our model in the later stages.
-Despite the recent emergence of consistent yet expensive scene calibration [44], video depth estimation [30] and motion tracking [53] estimators, one of our method’s main objectives is speed and simplicity.
-Thus, we focus on real-time out-of-the-box solutions to maintain a fast total reconstruction time (including prior computation).
-First, we start by applying sparse structure-from-motion (SfM) using COLMAP [40, 41] to estimate camera intrinsics and per-frame poses (extrinsics) $E=(e_{1},...,e_{N})$.
+To regularize the reconstruction from monocular video, previous approaches [GLT22, LNSW21, LGM23] apply estimates from several image processing applications, replacing the multi-view consistency signal with learned, data-driven scene analysis.
+Following this idea, we first extract additional priors, which we use to initialize and guide our model in the later stages.
+Despite the recent emergence of consistent yet expensive scene calibration [SCTS24], video depth estimation [LHS20] and motion tracking [WCC23] estimators, one of our method’s main objectives is speed and simplicity.
+Thus, we focus on fast out-of-the-box solutions to maintain a swift total reconstruction time (including prior computation).
+First, we start by applying sparse structure-from-motion (SfM) using COLMAP [SF16, SZFP16] or VGGSfM [WKRN24] to estimate camera intrinsics and per-frame poses (extrinsics) $E=(e_{1},...,e_{N})$.
 As a side product, we extract the sparse, globally-aligned SfM point cloud.
-We further use RAFT [47] to estimate dense optical flow in the forward and backward directions, and estimate per frame (dis-)occlusion masks based on forward-backward consistency.
-To reduce the depth ambiguity of dynamic scene content, we employ the fast monocular depth estimator DepthAnything [58], which provides relative disparity maps for every video frame.
-Instead of using a relative depth loss [25], we align the predicted disparity with the global SfM point cloud by estimating a scale and shift via RANSAC-based least squares and temporal smoothing.
+To reduce the depth ambiguity of dynamic scene content, we employ the fast monocular depth estimator DepthAnything [YKH24], which provides relative disparity maps for every video frame.
+Instead of using a relative depth loss [LNSW21], we align the predicted disparity with the global SfM point cloud by estimating a scale and shift via RANSAC-based least squares and temporal smoothing.
 This way, we obtain globally-aligned depth which can explicitly be integrated into our model.
+Alternatively, our method also accepts (sparse) metric LiDAR depth and camera poses obtained, e.g., via IMUs, which completely lifts the dependency on SfM calibration and point cloud priors.
 Lastly, we obtain binary segmentation masks $\omega$ of moving foreground object(s) to mask out dynamic scene content in our reconstruction and COLMAP SfM feature extraction.
 Similar to monocular depth, automatic motion segmentation is highly ill-posed, even in the presence of semantic segmentation masks.
-Thus, we use the semi-automatic video object segmentation system Cutie [10], which is guided by minimal user input in the form of manual keypoint selection.
+Thus, we use the semi-automatic video object segmentation system Cutie [COP24], which is guided by minimal user input.
 In line with our demands on simplicity, Cutie generates masks of sufficient quality in less than a minute, requiring as few as two mouse clicks, which in the future could be integrated into the data recording setup on mobile devices.
 3.2 Dynamic Neural Point Cloud Representation
-Our method represents the 4D scene as a Dynamic Neural Point Cloud, the first temporal extension of implicit neural point clouds [17].
-The primary idea of implicit point clouds is to replace the persistent point cloud $(p,f)$, consisting of a set of 3D positions $p$ and corresponding color or feature descriptor vectors $f$, with a sampling distribution $\mathcal{P}$ and a neural feature grid $\mathcal{A}$.
+Our method represents the 4D scene as a Dynamic Neural Point Cloud, the first temporal extension of implicit neural point clouds [HFK24].
+In their paper, Hahlbohm et al.
+propose a methodology that diverges from the conventional approach of storing both the spatial and photometric properties of the scene within a unified data structure.
+Instead, they advocate for a division of these properties and an implicit optimization of both.
+The geometric structure is depicted as an octree-based point probability field ($\mathcal{P}$), which undergoes progressive subdivision to ensure a uniform probability distribution across all leaf nodes.
+Conversely, the appearance features are embedded in an implicit coordinate-based multi-resolution hash grid ($\mathcal{A}$).
+During rendering, $\mathcal{P}$ is employed as an estimator for point positions, utilizing either random positions in each leaf or fixed sampling patterns.
+Meanwhile, per-point appearance features are queried from $\mathcal{A}$.
+The authors then employ fast bilinear splatting to render the resulting explicit point cloud, where gradients are backpropagated through their pipeline to the implicit representation [HFK24].
+Thus, the primary idea of implicit point clouds is to replace the persistent point cloud $(p,f)$, consisting of a set of 3D positions $p$ and corresponding color or feature descriptor vectors $f$, with a sampling distribution $\mathcal{P}$ and a neural feature grid $\mathcal{A}$.
 By repeatedly sampling and rendering an explicit point cloud $((p,\mathcal{A}(p)),p\sim\mathcal{P})$, it is possible to iteratively update and refine the spatial distribution, while maintaining all benefits of point-based rendering.
@@
 Given the natural imprecisions and relatively low resolution of monocular reconstruction, we find that the octree-based data structure suggested by Hahlbohm et al.
-[17] does not yield significant benefits due to the absence of fine geometric details.
+does not yield significant benefits due to the absence of fine geometric details.
 Instead, we use a fixed-resolution sparse voxel grid representation for the static and dynamic fields.
 Voxel Grid Initialization.
-Despite using a sparse data structure, starting from a fully occupied grid (i.e., all cells are sampled equally for all timestamps) is exceedingly inefficient regarding both memory consumption and model convergence.
+Despite using a sparse data structure, starting from a fully occupied grid (i.e., all cells are sampled equally for all timestamps) is extremely inefficient regarding both memory consumption and model convergence.
 Thus, we make use of the data-driven priors described in Sec.
@@
 Then, for every timestamp $t_{i}\in T$, we use the corresponding training camera matrix $e_{i}$ to project a set of random points in each voxel to the image plane, and gather the minimal relative distance to the estimated monocular depth $d_{i}$, and the maximum over the binary foreground segmentation mask $\omega_{i}$.
-We set the initial opacity-based sampling probability $\sigma_{d}(t_{i})$ of each cell to the relative deviation from the monocular depth, using the maximum over all timestamps for the static field ($\sigma_{s}=\max(\{\sigma_{d}(t)\>|\>t\in T\})$).
+For the dynamic probability field $\mathcal{P}_{d}$, we set the initial opacity-based sampling probability $\sigma_{d}(t_{i})$ of each cell based on the relative deviation ($d\in[0,1]$) from the monocular depth: $\sigma_{d}(t_{i})=(1-d(t_{i}))^{x}$ with $x$ being a free hyperparameter, to initially sample more points close to the expected surface.
+For the static probability field ($\mathcal{P}_{s}$), we instead use the maximum over all timestamps ($\sigma_{s}=\max\{\sigma_{d}(t_{i})\>|\>t_{i}\in T\}$).
 Additionally, we carve the temporal slices of $\mathcal{P}_{d}$ using the dynamic segmentation.
@@
 To obtain the final point positions $p$, we interpret the per-voxel opacity as a discrete PDF, using discrete multinomial sampling (with replacement) to propose a list of voxels, in each of which we place a single uniformly sampled 3D position.
-This way, more points are sampled in areas with a high expected scene density, which can iteratively be refined during optimization.
+This way, dense point clouds are sampled in clusters of high expected scene density, while empty or homogeneous static areas are sampled sparsely, enabling the renderer to infer depth ordering and occlusions via alpha-blending.
 For the resulting set of point positions, we further trace a binary mask $m_{p}$ indicating if a point was sampled from static or dynamic scene regions, enabling efficient model evaluation by omitting excessive queries of the dynamic appearance model.
 Voxel Grid Update.
-We update the sampling distribution field by tracking the blending weights assigned to every sampled explicit point during rasterization.
-We first apply a constant decay factor to the values of every unique voxel proposed by the multinomial sampling, and then assign it the maximum over the current value and the blending weights of all points sampled from the specific cell.
+We update the sampling distribution field by tracking the blending weights $(b_{i}\in\mathcal{B})$ assigned to every sampled explicit point during rasterization.
+We first apply a constant decay factor $(\gamma)$ to the values of every unique voxel proposed by the multinomial sampling, and then assign it the maximum over the current value and the blending weights of all points sampled from the specific cell:
+$\sigma_{s}=\max\{\sigma_{s}\gamma,\ max\{\mathcal{B}\}\},$ (1)
+where $\mathcal{B}$ is the set of blending weights, i.e., the visibility of all points sampled from this cell, which is calculated during alpha-blending at the rasterization step.
 For samples from dynamic voxels of $\mathcal{P}_{d}$, we apply the same procedure to the tracked static-to-dynamic ratios $\delta$ using per-point densities described in Sec.
 3.2.2.
-Afterwards, we prune voxels from the sampling distribution by threshold: For $\mathcal{P}_{d}$, we remove cells with a low $\delta$-value, as it either describes empty space or is well handled by the static model.
-We further remove voxels $\mathcal{P}_{s}$ from $\mathcal{P}_{d}$ whose expected opacity $\sigma$ is low in the static 3D field and for all remaining timestamps in the dynamic field.
+Afterwards, we prune voxels from the sampling distribution by threshold: For $\mathcal{P}_{d}$, we remove cells with a low sampling probability, as it either describes empty space or is well handled by the static model.
+We further remove static voxels from $\mathcal{P}_{s}$ whose expected opacity $\sigma$ is low both in the static 3D field and all timestamps in the dynamic field.
 This way, we progressively reduce memory consumption and refine importance sampling.
@@
 In a conventional point cloud, every point is assigned a fixed feature vector describing its appearance.
-As implicit neural point clouds [17] repeatedly extract new sets of point positions, they use a fast multi-resolution hash-encoded feature grid [33] $\mathcal{A}$ to model the appearance over the entire spatial sampling distribution.
+As implicit neural point clouds [HFK24] repeatedly extract new sets of point positions, they use a fast multi-resolution hash-encoded feature grid [MESK22] $\mathcal{A}$ to model the appearance over the entire spatial sampling distribution.
 To represent time-conditioned appearance for dynamic scenes, we again use a static 3D and dynamic 4D hash-encoded neural feature grid $\mathcal{A}=(\mathcal{A}_{s},\mathcal{A}_{d})$.
 As previously described by Wang et al.
-[51], using an ensemble of 3D and 4D hash-grids instead of a single 4D representation effectively improves the capabilities of the dynamic model by reducing hash collisions.
+[WCW23], using an ensemble of 3D and 4D hash-grids instead of a single 4D representation effectively improves the capabilities of the dynamic model by reducing hash collisions.
 To model the static and dynamic scene appearance at a spatial input position $p$ at time $t$, the appearance fields yield a density $\delta\in\mathbb{R}_{0}^{+}$ and an $n_{f}$-dimensional feature vector $f\in[0,1]^{n_{f}}$:
@@
 Note that we only have to query $\mathcal{A}_{d}$ for points indicated as dynamic by the binary segmentation mask $m$, which significantly reduces compute and memory.
-Following the formulation of D2NeRF [56], we then calculate the combined per-point density $\delta_{c}$ as the sum of static and dynamic densities, and obtain the final feature descriptor $f_{c}$ via density-weighted interpolation:
-$\delta_{c}=\begin{cases}\delta_{s}+\delta_{d}&\text{if}\quad m(p)\\ \delta_{s}&\text{else}\end{cases},\qquad f_{c}=\begin{cases}\frac{\delta_{s}f_{s}+\delta_{d}f_{d}}{\delta_{c}}&\text{if}\quad m(p)\\ f_{s}&\text{else}\end{cases}.$ (2)
+Following the formulation of D2NeRF [WZT22], we then calculate the combined per-point density $\delta_{c}$ as the sum of static and dynamic densities, and obtain the final feature descriptor $f_{c}$ via density-weighted interpolation:
+$\delta_{c}=\begin{cases}\delta_{s}+\delta_{d}&\text{if}\quad m(p)\\ \delta_{s}&\text{else}\end{cases},\qquad f_{c}=\begin{cases}\frac{\delta_{s}f_{s}+\delta_{d}f_{d}}{\delta_{c}}&\text{if}\quad m(p)\\ f_{s}&\text{else}\end{cases}$ (3)
 For rendering and probability field updates, we then transform density values $\delta$ into opacity $\sigma$, and calculate the dynamic ratio $\beta$ via:
@@
 3.3 Neural Point Cloud Rendering
-Given the time-dependent neural point cloud for an input view consisting of point positions $p$, point opacity $\sigma$, and appearance features $f$, we implement a fast differentiable point rasterizer to generate alpha ($\alpha$), depth ($d$), and 2D image feature maps $f_{\text{img}}$.
+Given the time-dependent neural point cloud for an input view consisting of point positions $p$, point opacity $\sigma$, and appearance features $f$, we implement a fast differentiable point rasterizer to generate alpha ($\alpha$), depth ($d$), and 2D image feature maps $f_{img}$.
 Naturally, the rasterized images still contain holes, a common characteristic of classical point rendering approaches.
-Thus, we use a neural renderer, implemented as a light-weight convolutional neural network with a U-Net [38] architecture to fill the remaining holes, and convert the neural feature map $f_{\text{img}}$ to a human-interpretable RGB image $c$.
+Thus, we use a neural renderer, implemented as a light-weight convolutional neural network with a U-Net [RFB15] architecture.
+This network translates the sparse rasterized feature image ($f_{img}$) to a dense output image ($c$), thus inferring human-interpretable RBG color from the higher dimension point cloud features while filling the remaining holes based on neighborhood information.
 In practice, we apply the rasterization and rendering components to the static, dynamic, and combined point cloud features separately, resulting in individual depth $d_{s,d,c}$ and color $c_{s,d,c}$ images that can be used for supervision and regularization.
@@
 The loss $\mathcal{L}_{\text{photo}}$ enforces photometric consistency between the rendered ($c$) and GT RGB image.
-Here, we use a weighted combination of pixel-level Cauchy loss [6], structural dissimilarity [54], and perceptual LPIPS [62] to maximize image quality.
+Here, we use a weighted combination of pixel-level Cauchy loss [BA96], structural dissimilarity [WBSS04], and perceptual LPIPS [ZIE18] to maximize image quality.
 Our depth regularization loss $\mathcal{L}_{\text{depth}}$ penalizes the difference between the rasterized dynamic depth $d_{d}$ and the aligned monocular depth prior $d_{\text{gt}}$ in the binary foreground mask $\omega$:
-$\mathcal{L}_{\text{depth}}=\frac{1}{\lvert\omega\rvert}\sum_{i=0}^{\lvert\omega\rvert}\omega_{i}\lvert d_{d,i}-d_{\text{gt},i}\rvert.$ (5)
-The segmentation regularizer $\mathcal{L}_{seg}$ encourages a clean separation between the static and dynamic models via skewed binary entropy [56] on the dynamic ratios $\beta$:
+$\mathcal{L}_{\text{depth}}=\frac{1}{\lvert\omega\rvert}\sum_{i=0}^{\lvert\omega\rvert}\omega_{i}\lvert d_{d,i}-d_{\text{gt},i}\rvert$ (6)
+The segmentation regularizer $\mathcal{L}_{seg}$ encourages a clean separation between the static and dynamic models via skewed binary entropy [WZT22] on the dynamic ratios $\beta$:
 $\mathcal{L}_{\text{seg}}=-\beta^{k}\log_{2}(\beta^{k})-(1-\beta^{k})\log_{2}(1-\beta^{k}),$ (7)
 where $k$ is the skewing bias.
-As our last regularizer $\mathcal{L}_{dist}$, we adapt the distortion loss of Mip-NeRF360 [4, 46] to encourage thin surfaces, as it is not feasible to reliably reconstruct the time-dependent volume of a deforming object from a single monocular observation.
-For a detailed description of our training and loss schedule, please see our supplemental material (Sec.
-A.2).
+As our last regularizer $\mathcal{L}_{dist}$, we adapt the distortion loss of Mip-NeRF360 [BMV22, SSC22] to encourage thin surfaces, as it is not feasible to reliably reconstruct the time-dependent volume of a deforming object from a single monocular observation.
+For a detailed description of our training and loss schedule, please see our supplemental material.
+Method GPU Hours [h] $\downarrow$ Frame Rate [FPS] $\uparrow$ Frame Rate [FPS] $\uparrow$
+(NVIDIA dataset) ($480{\times}270$) ($1920{\times}1080$)
+RoDynRF [LGM23] 28 ${<}1$
+HyperNeRF [PSH21] 64 ${<}1$
+DynamicNeRF [GSKH21] 74 ${<}0.1$
+NSFF [LNSW21] 223 ${<}1$
+DynIBaR [LWC23] 320 ${<}1$
+Marbles [SHU24] 3.5 200+
+CasualFVS [LZBM25] 0.25 48
+4D-GS [WYF24] 1.1 45 30
+Ours 0.27 78 22
+$\dagger$ Values gathered from literature – Values not reported
+Table 1: Performance Comparison. We report training times on the NVIDIA dataset, and inference frame rates for training- and FHD resolution.
+TiNeuVox RoDynRF 4D-GS Ours Ground Truth
+Figure 2: Qualitative comparisons on the NVIDIA dataset. Our method preserves fine details and clean foreground-background transitions.
+[…]
+Figure 3: Comparison on the iPhone dataset.
+We compare on five challenging scenes of the iPhone dataset.
+Our method better preserves the object shapes and fine details despite complex deformations.
+Methods indicated by $\dagger$ are the highest quality versions provided by Gao et al.
+[GLT22].
+The areas highlighted in red indicate lack of co-visibility between training and validation.
 4 Experimental Evaluation
@@
 4.1) and qualitative (Sec.
-4.2) evaluation and comparisons against state-of-the-art methods on two benchmark datasets: First, we evaluate our approach on the NVIDIA dynamic scenes dataset [59] (MIT license), featuring forward-facing scenes captured by a multi-view rig, monocularized by using a single frame per time to get a smooth camera trajectory.
-We extract training data using the procedure of RoDynRF [27], resulting in $12$ training and test views.
-For comparisons against DynIBaR [26] using their train/test split containing $\sim 150$ training views per scene, please see our supplemental material.
-Additionally, we evaluate our approach on the iPhone dataset [15] (Apache 2.0 license) which comprises challenging sequences with complex camera and scene motion trajectories, fine texture details, and (self-)occlusions.
-Furthermore, we provide an ablation study assessing the influence of individual pipeline components in Sec.
+4.2) evaluation and comparisons against state-of-the-art methods on two benchmark datasets: First, we evaluate our approach on the NVIDIA dynamic scenes dataset [SYKG20], featuring forward-facing scenes captured by a multi-view rig, monocularized by using a single frame per time to get a smooth camera trajectory.
+We extract training data using the procedure of RoDynRF [LGM23], resulting in $12$ training and test views.
+Additionally, we evaluate our approach on the iPhone dataset [GLT22] which comprises challenging sequences with complex camera and scene motion trajectories, fine texture details, and (self-)occlusions.
+We further show qualitative results including foreground-background separation on high-resolution (1080p) in-the-wild scenes from the DAVIS dataset [PPTM16], and provide an ablation study assessing the influence of individual pipeline components in Sec.
 4.3.
+For further experiments, including comparisons against DynIBaR [LWC23] and results on the HyperNeRF dataset [PSH21], please see our supplemental PDF.
 4.1 Quantitative Results
-Table 1: Timings on the NVIDIA dataset.
-We report training time and inference frame rates on the NVIDIA dataset with an image resolution of $480{\times}270$ pixels.
-Values for methods marked with $\dagger$ were gathered from the literature.
-Method GPU Hours [h] $\downarrow$ Frame Rate [FPS] $\uparrow$
-RoDynRF [27] 28 ${<}1$
-HyperNeRF [35] 64 ${<}1$
-DynamicNeRF [14] 74 ${<}0.1$
-NSFF [25] 223 ${<}1$
-DynIBaR [26] 320 ${<}1$
-4D-GS [55] 1.1 45
-Ours 0.27 70
+We compare the performance of our approach to several related benchmark methods.
+In addition to recent NeRF-based approaches, we also compare against the official implementation of 4D-GS [WYF24], a fast state-of-the-art forward rendering approach based on Gaussian splatting.
+For both 4D-GS and our method, we measure frame rates using the procedure described by SMERF [DHR24], rendering the test set $100$ times using a single RTX 4090 GPU.
+Please note that we can not report all numbers for concurrent (unpublished at time of submission) methods [LZBM25, WYG24, SHU24, LWH24] due to missing source code or differences in evaluation procedure (e.g., training resolution or camera calibration on the iPhone dataset).
+NVIDIA Average
+Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+T-NeRF [PCPMMN21] 18.33 0.436 0.511
+D-NeRF [PCPMMN21] 21.49 0.629 0.320
+HyperNeRF [PSH21] 17.60 0.377 0.502
+NR-NeRF [TTG21] 19.69 0.510 0.434
+NSFF [LNSW21] 24.33 0.751 0.257
+DynamicNeRF [GSKH21] 26.10 0.837 0.142
+RoDynRF [LGM23] 25.89 0.854 0.110
+TiNeuVox [FYW22] 19.74 0.501 0.412
+4D-GS [WYF24] 22.89 0.731 0.237
+Ours 25.64 0.845 0.109
+iPhone Average
+Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
+HyperNeRF+B+D+S [PSH21] 16.81 0.569 0.332
+Nerfies+B+D+S [PSB21] 16.45 0.570 0.339
+T-NeRF [GLT22] 15.70 0.538 0.458
+T-NeRF+B+D [GLT22] 17.06 0.573 0.390
+T-NeRF+B+D+S [GLT22] 16.96 0.577 0.379
+RoDynRF [LGM23] 17.00 0.535 0.517
+Marbles [SHU24] 16.72 0.418
+4D-GS [WYF24] 13.87 0.459 0.407
+Ours 16.04 0.589 0.348
+[…]
 Table 2: Quantitative comparisons on the NVIDIA and iPhone datasets against recent dynamic view synthesis approaches.
@@
 The three best results are highlighted in green in descending order of saturation.
-NVIDIA Average
-Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-T-NeRF [37] 18.33 0.436 0.511
-D-NeRF [37] 21.49 0.629 0.320
-HyperNeRF [35] 17.60 0.377 0.502
-NR-NeRF [49] 19.69 0.510 0.434
-NSFF [25] 24.33 0.751 0.257
-DynamicNeRF [14] 26.10 0.837 0.142
-RoDynRF [27] 25.89 0.854 0.110
-TiNeuVox [11] 19.74 0.501 0.412
-4D-GS [55] 22.89 0.731 0.237
-Ours 25.64 0.845 0.109
-iPhone Average
-Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
-HyperNeRF+B+D+S [35] 16.81 0.569 0.332
-Nerfies+B+D+S [34] 16.45 0.570 0.339
-T-NeRF [15] 15.70 0.538 0.458
-T-NeRF+B+D [15] 17.06 0.573 0.390
-T-NeRF+B+D+S [15] 16.96 0.577 0.379
-RoDynRF [27] 17.00 0.535 0.517
-4D-GS [55] 13.87 0.459 0.407
-Ours 15.79 0.585 0.354
-[…]
-We compare the performance of several related methods in Tab.
-1.
-For all NeRF-based approaches, we report values from the literature.
-To further compare against a fast forward rendering approach, we train the state-of-the-art Gaussian splatting-based 4D-GS [55] on both benchmark datasets.
-For both 4D-GS and our method, we measure performance on a single RTX 4090 GPU.
-As evident from the results, our method is significantly faster than backward rendering approaches, and even trains faster than 4D-GS while achieving similar frame rates during inference.
+As evident from Tab.
+1, our method is significantly faster than backward rendering approaches, and even trains faster than 4D-GS while achieving similar frame rates during inference.
 A breakdown of timings for individual method components is provided in the supplemental material.
-We further compare image quality based on PSNR, SSIM [54] and LPIPS [62] quality metrics on the NVIDIA and iPhone datasets in Tab.
+We further compare image quality based on PSNR, SSIM [WBSS04], and LPIPS [ZIE18] quality metrics on the NVIDIA and iPhone datasets in Tab.
 2.
 To avoid inconsistencies across the reported values, we use the official results and models kindly provided by Liu et al.
-[27], Li et al.
-[26], and Gao et al.
-[15] to consistently recalculate all metrics using the VGG backbone for LPIPS calculation.
-For the iPhone dataset, we use masked metric calculation (mPSNR, mSSIM, mLPIPS) as suggested and implemented by Gao et al.
-[15] to account for co-visibility.
+[LGM23], Li et al.
+[LWC23], and Gao et al.
+[GLT22] to consistently recalculate all metrics, using the VGG backbone for LPIPS calculation.
+On the iPhone dataset, we instead use masked metric calculation (mPSNR, mSSIM, mLPIPS) as suggested and implemented by Gao et al.
+[GLT22] to account for co-visibility.
+For our method, we further report values for two versions, trained with our monocular depth alignment (Ours) and the provided ${}_{\text{Lidar}}$).
 Overall, we find our method is not only the fastest in terms of reconstruction and rendering speed but also achieves competitive image quality, especially in terms of perceptual image quality metrics (LPIPS).
@@
 4.2 Qualitative Results
-TiNeuVox RoDynRF 4D-GS Ours Ground Truth
-Figure 2: Qualitative comparisons on the NVIDIA dataset. Our method preserves fine details and clean foreground-background transitions.
-[…]
-Figure 3: Comparison on the iPhone dataset.
-We compare on two challenging scenes of the iPhone dataset.
-Our method better preserves the object shapes and fine details despite complex deformations.
-Methods indicated by $\dagger$ are the highest quality versions provided by Gao et al.
-[15]
-We show qualitative results on the NVIDIA (Fig.
+We show qualitative comparisons on the NVIDIA (Fig.
 2) and iPhone (Fig.
 3) datasets.
-For comparisons on the iPhone sequences, we show the best version of HyperNeRF [35] and T-NeRF [37] according to image quality metrics, using all additional regularizers (Lidar depth, random background, and distortion loss) proposed by Gao et al.
-[15].
-For our method, we report values for two versions, trained with the provided Lidar or monocular depth data.
+For the iPhone sequences, we show the best version of HyperNeRF [PSH21] and T-NeRF [PCPMMN21] according to image quality metrics, using all additional regularizers (Lidar depth, random background, and distortion loss) proposed by Gao et al.
+[GLT22].
 Again, we find that our method achieves a competitive and oftentimes superior perceived image quality, preserving higher frequencies and clean edges.
-For dynamic view synthesis results, please see our supplemental video.
+As our 4D model is temporally conditioned and does not directly estimate inter-frame motion, it can well handle scene movements of varying complexity (e.g., the rotating motion on the Paper-windmill scene in Fig.
+3 or human motion such as in the Skating scene in LABEL:{fig:nvidia}), as long as it is constantly observed from the input camera trajectory (and not occluded).
+Figure 4: Novel views synthesized using complex in-the-wild sequences of the DAVIS dataset. On the right, we show the corresponding visualizations of our implicit foreground/background separation.
+Figure 4 further provides novel views synthesized using complex in-the-wild sequences from the DAVIS dataset.
+We additionally include visualizations of our implicit foreground/background separation obtained by only rendering the plain static or dynamic components of the sampled point clouds.
+Given that our method is efficient in both memory and compute, we train our DAVIS models on the full original resolution (FHD) and double the sampled point cloud size to $8M$ points to maintain a dense distribution in the rasterized feature images.
+This upgrade in rendering resolution and point queries increases the training time to $0.52$ hours and reduces online framerates to $18$ fps.
+Overall, our experiments show that DNPC robustly handles high-resolution in-the-wild inputs provided that decent camera poses can be obtained from current SfM methods.
+For more video results, including foreground/background separation, 3D anaglyphs renderings and live GUI capture, please see our supplemental material.
 4.3 Ablation Study
-Table 3: Comparison of ablated D-NPC versions in terms of image quality, model size and speed.
 Method Image Quality Model Size Frame Rate
@@
 Oursw/o U-Net 24.24 0.798 0.155 662 78
+Table 3: Comparison of ablated D-NPC versions in terms of image quality, model size, and speed.
 To assess the improvements enabled by our main pipeline components, we optimize three variations of our method on the NVIDIA dataset, and compare them to our full D-NPC model (full).
@@
 5 Discussion and Limitations
-Figure 4: Failure Cases. Top: Cropped foreground not contained in the training camera frustum. Bottom: Object duplications due to depth misalignment.
-In our experiments, our D-NPC yields the best trade-off between optimization time and image quality and is currently the only monocular method providing real-time frame rates for interactive applications.
-Especially in terms of perceptual LPIPS and high image frequencies, our method quantitatively and visually outperforms recent monocular view synthesis methods while only requiring a fraction of the optimization time.
+The conducted experiments demonstrate the adaptability and effectiveness of the implicit neural point cloud representation [HFK24] for non-rigid novel view synthesis; Our temporal extension D-NPC effectively leverages data-driven priors for dynamic reconstruction from single monocular video, yielding a favorable trade-off between optimization time and image quality.
+Due to the fast forward-rendering approach, it is currently one of the only purely monocular methods enabling real-time frame rates for interactive scene exploration.
+Especially in terms of perceptual LPIPS and high image frequencies, our method quantitatively and visually outperforms most recent monocular view synthesis methods while only requiring a fraction of the optimization time.
 However, there is room for improvement: Our method relies on a decent quality of camera poses and monocular depth.
 Moreover, our initialization process assumes that the dynamic content is fully contained in the camera frustum, resulting in visible crops when synthesizing far-off novel views, as shown in Fig.
-4.
+5.
+Finally, our method handles videos of up to $\approx 500$ frames (similar to related approaches) using less than $24$GB of GPU-VRAM at FHD resolution, with longer video sequences remaining a limitation.
+Figure 5: Failure Cases. Top: Cropped foreground not contained in the training camera frustum. Bottom: Object duplications due to depth misalignment.
 6 Conclusion
-We presented Dynamic Neural Point Clouds, an efficient implicit point cloud-driven [17] approach for non-rigid novel-view synthesis from monocular video inputs.
+We presented Dynamic Neural Point Clouds, an efficient implicit point cloud-driven [HFK24] approach for non-rigid novel-view synthesis from single monocular video inputs.
 In contrast to existing ray marching-based NeRF methods, our rasterization-based forward rendering approach enables a direct and elegant way to initialize and guide the reconstruction from data-driven priors, not only making it one of the fastest approaches in terms of optimization time, but also enabling rendering at real-time frame rates.
 While our experiments show our novel approach already yields competitive image quality according to several metrics on common benchmark datasets, we believe that our work constitutes a promising new direction for future research and extensions, like the integration of diffusion-based priors.
-Acknowledgments and Disclosure of Funding
+Acknowledgments
 We thank Jann-Ole Henningson for help with figures and writing.
-This work was partially funded by the German Research Foundation (DFG) projects “Immersive Digital Reality” (ID 283369270) and “Real-Action VR” (ID 523421583).
-Appendix A Appendix – Supplemental Material
-In this supplemental, we provide details on our model configuration (Sec.
-A.1) and optimization procedure (Sec.
-A.2), include additional results (Sec.
-A.3), and discuss concurrent work on arXiv (Sec.
-A.4) as well as the societal impact (Sec.
-A.5) of our work.
-A.1 Model Details
-Table 4: Appearance Field Hash Grid Configurations.
-Parameter 3D Static Grid 4D Dynamic Grid
-Base Resolution 16 16
-Resolution Levels 10 8
-Per Level Scale 2 2
-Features Per Level 4 4
-Logarithmic Size 21 22
-In our model, the crucial neural components regarding the image quality are the appearance feature grids and the neural rendering network.
-To model our static and dynamic appearance fields, we use the fast multi-resolution hash encoding implementation by Müller et al.
-[33, 32] After initially carving our sampling distribution voxel grid using the monocular depth and dynamic segmentation priors, we set the spatial scale of our 3D and 4D feature grids to the minimum axis-aligned bounding box enclosing all remaining static and dynamic voxels, respectively.
-We parameterize the hash-grids according to the values in Tab.
-4, giving more total capacity to the higher dimensional dynamic representation.
-Both hash encodings are followed by a shallow ReLU-activated MLP with one hidden $64$-neuron layer to resolve hash collisions.
-For our neural rendering network used for hole-filling, we apply a simple image-to-image translation approach based on a U-Net architecture [38].
-In contrast to Hahlbohm et al.
-[17], we do not use any multi-resolution inputs.
-Instead, we use two downsampling and consecutive upsampling blocks, each consisting of two ReLU-activated 2D convolutions with kernel size $3$, followed by average pooling / bilinear upsampling.
-During down- and upscaling, each convolution block doubles/halves the amount of feature channels while adjusting the spatial resolution, going from an initial $64$ channels at the finest to a maximum of $256$ channels at the coarsest spatial resolution.
-A.2 Optimization Details
-Table 5: Loss Hyperparameters.
-Hyperparameter Value
-$\lambda_{\text{photo, Cauchy}}$ 1.0
-$\lambda_{\text{photo, D-SSIM}}$ 0.9
-$\lambda_{\text{photo, LPIPS}}$ 0.05
-$\lambda_{\text{depth}}$ 20.0
-$\lambda_{\text{seg}}$ $1\mathrm{e}{-3}$
-$\lambda_{\text{dist}}$ 1.0
-$k$ 0.5
-We train our model for a total of 10K iterations on a consumer-grade RTX 4090 GPU.
-During each iteration, we sample an explicit point cloud consisting of $4$M points (i.e., a batch of spatial 3D locations).
-While the sampling distribution is manually updated based on statistics of the extracted explicit point clouds (see Sec.
-3.2.1), the differentiable rasterizer and neural rendering network enable end-to-end optimization using gradient descent for the network and hash-grid parameters.
-We use the Adam [20] optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1\mathrm{e}{-15}$) with an initial learning rate of $1\mathrm{e}{-2}$ for the appearance hash-grids and $5\mathrm{e}{-4}$ for the rendering network, which are exponentially decayed to $\frac{1}{3}\mathrm{e}{-3}$ and $5\mathrm{e}{-5}$ during optimization.
-For our loss hyperparameters (Sec.
-3.4), we use the values provided in Tab.
-5.
-Note that, after initially training with a high value $\lambda_{depth}$ to guide the optimization in a reasonable direction, we turn off our depth loss after $500$ training iterations to allow the model to recover from errors in the monocular depth estimates.
-In contrast to D2NeRF, who use $k=2$ to bias the binary segmentation loss towards static background, which usually constitutes a majority of the scene, we use a value of $0.5$ to favor the dynamic model.
-This is due to the fact that our initialization procedure already discards large amounts of dynamic model evaluation in static areas.
-As a result, the remaining dynamic samples are already biased towards potentially dynamic areas, which we reflect in the choice of skew $k$.
-In addition to the losses already discussed in Sec.
-3.4, we further apply weight decay (i.e., an L2 loss) with a factor of $2.5$ on the appearance grid parameters [5].
-In practice, we find that this decay helps to prevent overfitting and improves the quality of generated novel views.
-Our Python / CUDA implementation will be made publicly available.
-A.3 Additional Results
-We provide detailed breakdowns on the NVIDIA and iPhone datasets in Tab.
-6 and Tab.
-7, respectively.
-As evident from the per-scene results, our method consistently achieves competitive results across a variety of challenging scenes, and particularly excels at perceptual LPIPS scores due to its ability to retain high-frequency details.
-Table 6: Quantitative comparison on the NVIDIA dataset.
-We compare our method against recent dynamic view synthesis approaches on the NVIDIA dynamic scenes dataset.
-Our method achieves the best scores in terms of perceptual image quality.
-The three best results are highlighted in green in descending order of saturation.
-Balloon1 Balloon2 Jumping Playground
-Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-T-NeRF [37] 18.54 0.448 0.416 20.69 0.573 0.335 18.04 0.496 0.541 14.68 0.240 0.527
-D-NeRF [37] 19.06 0.492 0.382 20.76 0.557 0.367 22.36 0.712 0.318 20.18 0.670 0.250
-HyperNeRF [35] 13.96 0.214 0.677 16.57 0.286 0.544 18.34 0.521 0.460 13.17 0.179 0.603
-NR-NeRF [49] 17.39 0.369 0.497 22.41 0.703 0.314 20.09 0.644 0.418 15.06 0.248 0.450
-NSFF [25] 21.96 0.701 0.281 24.27 0.731 0.276 24.65 0.813 0.227 21.22 0.705 0.263
-DynamicNeRF [14] 22.36 0.775 0.169 27.06 0.859 0.113 24.68 0.842 0.153 24.15 0.849 0.145
-RoDynRF [27] 22.37 0.782 0.157 26.19 0.846 0.107 25.66 0.853 0.129 24.96 0.899 0.080
-TiNeuVox [11] 17.30 0.392 0.500 19.06 0.458 0.396 20.81 0.641 0.399 13.84 0.213 0.541
-4D-GS [55] 20.82 0.666 0.271 24.34 0.775 0.211 22.16 0.735 0.270 20.22 0.724 0.192
-Ours 22.26 0.779 0.161 25.84 0.835 0.107 24.51 0.839 0.116 23.59 0.866 0.099
-Skating Truck Umbrella Average
-Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-T-NeRF [37] 20.32 0.551 0.562 18.33 0.462 0.523 17.69 0.284 0.670 18.33 0.436 0.511
-D-NeRF [37] 22.48 0.678 0.403 24.10 0.743 0.213 21.47 0.554 0.310 21.49 0.629 0.320
-HyperNeRF [35] 21.97 0.640 0.322 20.61 0.493 0.394 18.59 0.308 0.517 17.60 0.377 0.502
-NR-NeRF [49] 23.95 0.745 0.342 19.33 0.485 0.548 19.63 0.376 0.471 19.69 0.510 0.434
-NSFF [25] 29.29 0.889 0.186 25.96 0.775 0.248 22.97 0.644 0.315 24.33 0.751 0.257
-DynamicNeRF [14] 32.66 0.951 0.077 28.56 0.872 0.146 23.26 0.711 0.190 26.10 0.837 0.142
-RoDynRF [27] 28.68 0.939 0.076 29.13 0.900 0.089 24.26 0.757 0.135 25.89 0.854 0.110
-TiNeuVox [11] 23.32 0.723 0.283 23.86 0.662 0.328 20.00 0.416 0.439 19.74 0.501 0.412
-4D-GS [55] 26.04 0.876 0.171 25.08 0.753 0.233 21.59 0.588 0.309 22.89 0.731 0.237
-Ours 30.22 0.948 0.061 28.92 0.897 0.084 24.15 0.747 0.136 25.64 0.845 0.109
-Table 7: Quantitative evaluation on the iPhone dataset.
-We compare our method against related approaches using masked metrics based on co-visibility [15].
-Methods marked with $\dagger$ are trained using the code and additional regularizers introduced by Gao et al.
-[15].
-We further show values for our method trained on monocular depth instead of the provided Lidar data.
-The three best results are highlighted in green in descending order of saturation.
-Apple Block Paper-windmill Space-out
-Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
-HyperNeRF+B+D+S [35] 17.64 0.743 0.478 17.54 0.670 0.331 17.38 0.382 0.209 17.93 0.605 0.320
-Nerfies+B+D+S [34] 17.54 0.750 0.478 16.61 0.639 0.389 17.34 0.378 0.211 17.79 0.622 0.303
-T-NeRF [15] 16.64 0.719 0.559 16.36 0.637 0.434 15.29 0.278 0.447 17.17 0.602 0.397
-T-NeRF+B+D [15] 17.32 0.720 0.514 17.60 0.670 0.375 17.56 0.351 0.295 18.34 0.619 0.345
-T-NeRF+B+D+S [15] 17.43 0.728 0.508 17.52 0.669 0.346 17.55 0.367 0.258 17.71 0.591 0.377
-RoDynRF [27] 18.74 0.723 0.552 18.05 0.634 0.513 16.71 0.321 0.482 18.57 0.594 0.413
-4D-GS [55] 15.35 0.686 0.432 13.40 0.544 0.506 14.48 0.201 0.336 14.37 0.513 0.366
-Ours 16.45 0.749 0.487 15.32 0.626 0.363 18.03 0.434 0.209 17.17 0.624 0.288
-[…]
-Spin Teddy Wheel Average
-Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
-HyperNeRF+B+D+S [35] 19.20 0.561 0.325 13.97 0.568 0.350 13.99 0.455 0.310 16.81 0.569 0.332
-Nerfies+B+D+S [34] 18.38 0.585 0.309 13.65 0.557 0.372 13.82 0.458 0.310 16.45 0.570 0.339
-T-NeRF [15] 16.36 0.504 0.530 13.38 0.550 0.475 14.73 0.477 0.362 15.70 0.538 0.458
-T-NeRF+B+D [15] 18.95 0.541 0.468 13.99 0.573 0.435 15.69 0.541 0.301 17.06 0.573 0.390
-T-NeRF+B+D+S [15] 19.16 0.567 0.443 13.71 0.570 0.429 15.65 0.548 0.292 16.96 0.577 0.379
-RoDynRF [27] 17.41 0.484 0.570 14.34 0.537 0.613 15.20 0.449 0.478 17.00 0.535 0.517
-4D-GS [55] 15.36 0.418 0.310 12.36 0.506 0.474 11.78 0.343 0.427 13.87 0.459 0.407
-Ours 17.07 0.567 0.322 12.63 0.541 0.475 13.85 0.555 0.330 15.79 0.585 0.354
-[…]
-Table 8: Foreground analysis on the short NVIDIA dataset.
-We show additional quantitative analysis on the short NVIDIA sequences.
-We use masked metrics [15] with the provided foreground masks to evaluate image quality on only the dynamic part of the scenes.
-The three best results are highlighted in green in descending order of saturation.
-Balloon1 Balloon2 Jumping Playground
-Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
-T-NeRF [37] 17.90 0.897 0.245 19.89 0.953 0.133 15.24 0.914 0.285 14.81 0.957 0.267
-D-NeRF [37] 18.75 0.907 0.294 18.23 0.950 0.203 19.23 0.941 0.160 16.65 0.970 0.227
-HyperNeRF [35] 14.43 0.865 0.471 15.95 0.931 0.251 14.82 0.913 0.278 12.32 0.950 0.390
-NR-NeRF [49] 15.87 0.880 0.383 20.13 0.962 0.177 16.76 0.934 0.228 13.91 0.956 0.311
-NSFF [25] 19.34 0.925 0.225 22.63 0.969 0.155 20.42 0.955 0.152 20.22 0.983 0.162
-DynamicNeRF [14] 19.73 0.931 0.134 23.79 0.981 0.042 20.68 0.958 0.100 21.05 0.984 0.095
-RoDynRF [27] 19.70 0.928 0.147 23.05 0.977 0.046 21.54 0.962 0.093 20.29 0.984 0.097
-TiNeuVox [11] 16.80 0.888 0.330 19.59 0.955 0.147 17.34 0.932 0.196 13.02 0.950 0.363
-4D-GS [55] 18.72 0.911 0.247 22.21 0.973 0.066 18.60 0.938 0.173 17.23 0.976 0.156
-Ours 19.76 0.930 0.159 22.14 0.973 0.056 21.07 0.960 0.080 19.76 0.982 0.107
-Skating Truck Umbrella Average
-Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
-T-NeRF [37] 22.56 0.993 0.086 18.40 0.913 0.409 16.32 0.945 0.496 17.87 0.939 0.274
-D-NeRF [37] 22.05 0.993 0.090 19.84 0.923 0.312 17.34 0.953 0.365 18.87 0.948 0.236
-HyperNeRF [35] 23.86 0.993 0.074 19.40 0.918 0.281 14.76 0.940 0.489 16.51 0.930 0.319
-NR-NeRF [49] 27.17 0.995 0.052 18.25 0.912 0.534 16.23 0.948 0.404 18.33 0.941 0.298
-NSFF [25] 31.37 0.998 0.019 26.51 0.971 0.135 19.28 0.967 0.238 22.82 0.967 0.155
-DynamicNeRF [14] 37.22 0.999 0.010 26.75 0.973 0.067 19.66 0.969 0.152 24.13 0.971 0.086
-RoDynRF [27] 34.37 0.997 0.035 27.29 0.978 0.061 20.85 0.970 0.134 23.87 0.971 0.087
-TiNeuVox [11] 27.18 0.996 0.047 21.39 0.934 0.204 16.09 0.949 0.358 18.77 0.943 0.235
-4D-GS [55] 30.99 0.997 0.048 22.77 0.947 0.207 18.36 0.956 0.259 21.27 0.957 0.165
-Ours 36.38 0.999 0.014 27.16 0.975 0.063 21.71 0.975 0.087 24.00 0.971 0.081
-As most natural scenes are dominated by static background, we individually assess the quality of dynamic foreground reconstruction, which is the most complex part of dynamic view synthesis from monocular inputs.
-To this end, we recalculate metrics on the NVIDIA sequences using masked metrics [15], based on the motion masks contained in the dataset.
-As shown in Tab.
-8, our method still performs favorably regarding perceptual metrics, indicating that improvements were not exclusively located in the static background.
-As our method can handle significantly longer and temporally denser sequences than the commonly used train/test splits of the NVIDIA dynamic scenes dataset, we further compare against DynIBaR [26], using their evaluation protocol containing $\sim 150$ training views per scene.
-For fairness, we also use the consistent video depth made available by the authors.
-Results are shown in  Tab.
-9.
-Table 9: Comparison on the long NVIDIA dataset.
-We compare to the state-of-the-art approach DynIBaR [26] on the long sequences from the NVIDIA dynamic scenes dataset.
-Masked metrics are calculated with respect to the dynamic foreground masks.
-Method Balloon1 Balloon2 Jumping Playground Skating Truck Umbrella DynamicFace Average
-PSNR $\uparrow$
-DynIBaR [26] 28.99 28.90 23.74 28.32 31.50 33.75 28.25 29.18 29.08
-Ours 27.32 30.09 22.67 25.51 33.29 32.14 24.76 28.18 28.00
-SSIM $\uparrow$
-DynIBaR [26] 0.920 0.942 0.866 0.953 0.969 0.966 0.898 0.971 0.936
-Ours 0.902 0.934 0.841 0.895 0.969 0.952 0.737 0.962 0.899
-LPIPS $\downarrow$
-DynIBaR [26] 0.064 0.070 0.126 0.049 0.058 0.049 0.085 0.052 0.069
-Ours 0.065 0.052 0.128 0.072 0.044 0.045 0.142 0.048 0.074
-mPSNR $\uparrow$
-DynIBaR [26] 23.91 25.53 17.49 20.91 19.52 29.06 25.96 30.55 24.12
-Ours 22.35 23.51 16.30 18.30 19.89 25.57 21.75 23.94 21.45
-mSSIM $\uparrow$
-DynIBaR [26] 0.944 0.985 0.911 0.989 0.990 0.988 0.971 0.997 0.972
-Ours 0.937 0.980 0.896 0.984 0.992 0.979 0.921 0.991 0.960
-mLPIPS $\downarrow$
-DynIBaR [26] 0.079 0.032 0.168 0.062 0.070 0.038 0.043 0.010 0.062
-Ours 0.077 0.037 0.193 0.097 0.069 0.053 0.110 0.026 0.083
-While our method does not quite attain the image quality of DynIBaR, we achieve competitive results with significantly lower training and rendering times.
-Table 10: Timings for individual method components.
-      Component       Time [s]
-      Data Preprocessing       140
-      Depth Alignment       6
-      Model Initialization       1
-      Model Optimization       817
-We further provide a breakdown of timings for individual components of our method on the NVIDIA dataset in Tab.
-10.
-Note that the cost of data preprocessing highly depends on the amount of input images, while the optimization and rendering times increase with image resolution.
-A.4 Concurrent Work
-Concurrent to our work, other methods on arXiv address dynamic reconstruction and novel-view synthesis from monocular video.
-Lee et al.
-[23] propose a new scene representation for fast reconstruction and interactive rendering from in-the-wild monocular videos.
-Similar to our approach, they separately model the foreground and background, but use plane based proxy geometry for static regions, and point clouds for dynamic foreground, enabling fast forward rendering and an impressive image quality.
-More details and results are presented at the official project page: https://casual-fvs.github.io.
-Following recent trends in static scene reconstruction, Wang et al.
-[50] integrate diffusion-based priors in the monocular reconstruction.
-While not being exceptionally fast, their approach hallucinates additional training supervision for occluded areas, improving results on the challenging iPhone dataset [15].
-Again, we refer the interested reader to the official project page for more information: https://mightychaos.github.io/dpdy_proj/.
-A.5 Societal Impact
-Our method reconstructs a 4D representation and generates novel views based on a monocular input video.
-As the method itself cannot fill unseen areas or use/provide any information that goes beyond the inputs provided by the users themselves, we do not expect any negative societal impact resulting from our work.
-In terms of environmental impact, the short training times and consumer-grade GPU (RTX 3090/4090) required by our method makes it very economic in terms of power consumption and hardware requirements: Overall, after downloading and setting up our code, running a single reconstruction, including data preprocessing with user inputs, takes under one hour of time.
-When running on a slightly above average PSU with $750$W at full capacity (which is usually not the case), given an average energy price of \$$0.173$ per kWh (April 2024 in the U.S.), we get a worst-case cost of \$$0.13$ per scene reconstruction.
-Thus, in comparison to DynIBaR [26], the current the state-of-the-art in terms of image quality, which requires $384$ GPU hours on high-end A100 GPUs, our method has a significantly lower environmental footprint.
+This work was partially funded by the DFG under projects “Real-Action VR” (ID 523421583), and “Increasing Realism of Omnidirectional Videos in Virtual Reality” (ID 491805996); the L3S Research Center, Hanover, Germany; as well as by the ERC Consolidator Grant 4DRepLy (770784).
+Open Access funding enabled and organized by Projekt DEAL.